### v1

# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

Abstract
Minkowski tensors, also known as tensor valuations, provide robust $n$-point information for a wide range of random spatial structures. Local estimators for voxelized data, however, are unavoidably biased even in the limit of infinitely high resolution. Here, we substantially improve a recently proposed, asymptotically unbiased algorithm to estimate Minkowski tensors for voxelized data. Our improved algorithm is more robust and efficient. Moreover we generalize the theoretical foundations for an asymptotically bias-free estimation of the interfacial tensors to the case of finite unions of compact sets with positive reach, which is relevant for many applications like rough surfaces or composite materials. As a realistic test case, we consider, among others, random (beta) polytopes. We first derive explicit expressions of the expected Minkowski tensors, which we then compare to our simulation results. We obtain precise estimates with relative errors of a few percent for practically relevant resolutions. Finally, we apply our methods to real data of metallic grains and nanorough surfaces, and we provide an open-source python package, which works in any dimension.
Keywords: tensor valuations, microstructure characterization, anisotropy, digitized image analysis
2020 MSC: 94A08 · 68U10 · 60D05 · 53C65 · 28A75 · 62H35 · 52A22
1 Introduction
Random spatial structures appear ubiquitously in nature and technology. Examples (and their corresponding mathematical models) are heterogeneous materials and porous media (random sets) [44, 64, 45, 1, 3], cellular tissues and foam-like structures (tessellations) [46, 50, 34, 60], rough surfaces (random fields) [2, 65, 59, 51], and bacterial colonies (particle processes) [12, 19, 66] as well as error-correcting code (high-dimensional sphere packings) [67, 13]. This variety of examples corresponds to a similar diversity in random shapes.
In all of these cases, Minkowski tensors (also known as tensor valuations) [55, 28] provide a comprehensive, unified shape analysis. The Minkowski tensors [57] are generalizations of the well-known Minkowski functionals (or intrinsic volumes) from integral geometry [55]; see Section 2 for more details and the theoretical background. The intrinsic volumes represent intuitive geometric information, e.g., on the volume, surface area, or curvature; the Minkowski tensors additionally distinguish different orientations with respect to these geometrical properties [37].
Minkowski tensors of different ranks capture symmetries of all orders. They quantify the degree of anisotropy as well as the preferred directions. They are comprehensive in the sense of Hadwiger’s and Alesker’s characterization theorems [55], intuitively speaking, they contain all motion-covariant, continuous, and additive shape information. Since they are additive, Minkowski functionals and tensors provide robust access to $n$-point information [41]. Moreover, the definition in real space allows for a convenient treatment of boundary conditions [56].
The versatility of Minkowski functionals and tensors is demonstrated by their ample successful applications in a wide range of fields, including statistical physics [40, 42, 33], biology [8, 7, 6], spatial statistics [66, 16], image analysis [44, 45], as well as astronomy and cosmology [18, 52, 30, 35, 15, 14]. Explicit examples of random spatial structures characterized via Minkowski functionals and tensors are porous media [5, 4, 36, 3, 29], composite materials [17], nanorough surfaces [59], fluid demixing [10], and trabecular bone [50, 37, 11].
A challenge, however, is the application of Minkowski tensors to voxelized data because the discrete set of available orientations introduces a systematic bias usually even in the limit of infinitely high resolution. For gray-scale data, marching-cube algorithms can turn the voxelized data into triangulated meshes with continuous orientations, which results in asymptotically bias-free algorithms [39, 38, 61, 62]. If no gray-scale data is available, any local algorithm on black-and-white voxels will be asymptotically biased, i.e., the estimator will not converge to the true value in the limit of infinite resolution [63].
An algorithmic solution to the problem was recently proposed in [21], which had been inspired by [43]. The key idea is to construct Voronoi cells that can encode global information and thus guarantee asymptotically bias-free estimators of all Minkowski tensors for sets of positive reach. This requirement of a positive reach, however, excludes many interesting applications, like rough surfaces and composite materials. More importantly, the proposed algorithm suffered from numerical instabilities since it relies on a matrix inversion.
Here, we substantially advance the robustness of the algorithm by replacing the matrix inversion by a least squares fit. We, moreover, increase its efficiency by using an unbiased estimator of the Voronoi volume integrals. This revised algorithm is also asymptotically unbiased for sets of positive reach, as we prove rigorously. We demonstrate its robustness and accuracy for exemplary cases of geometric shapes, stochastic models, and real data.
Even for a simple example of a set without positive reach, our algorithm achieves surprisingly good results (with relative deviations of less than 3% in the eigenvalues of the interfacial tensors, see Table 6). Since such precision is not theoretically substantiated, we also introduce an alternative method to estimate interfacial tensors, for which we are able to prove that it is asymptotically unbiased even for finite unions of compact sets with positive reach.
Together with the paper, we publish an open-source python package including both methods. It is the first implementation of Minkowski tensors that we know of that can be applied to any dimension. It can also be applied to any discrete representation of the set, i.e., not only for pixelated images, but representations using non-cubic lattice or even random point patterns. The last generalization is especially helpful in high dimensions, where similar to numerical integration, lattice representations may converge too slowly. Details on the algorithm, implementation, and suitable choices of parameters are discussed in Sections 4.14.3.
We demonstrate the reliability of our methods for simple geometric test cases in Sections 4.4 and 4.5. For a stochastic test case, we apply our estimator to a classic example of random convex sets, isotropic random polytopes (see Section 5), both in theory (see Section 5.1) and simulations (see Section 5.2).
In Section 6, we turn to experimental data. First, we analyze metallic grains in a polycrystalline nickel-based superalloy using data from [60]. The Minkowski tensors characterize the cells to be distinctly more anisotropic with respect to curvature than the surface area (see Section 6.1). Then, we determine the interfacial tensors of nanorough surfaces from [59] (see Section 6.2).
2 Minkowski tensors in integral geometry
[…]
  • They form a basis of the vector space of continuous, isometry invariant, additive, functions on ${\mathcal{K}}^{d}$; $V_{i}$ is positively homogeneous of degree $i$ [26, Theorem 4.20].
  • They arise as coefficient functionals of a Steiner formula [26, Theorem 3.10].
  • They satisfy various integral geometric formulas such as Crofton formulas, projection formulas and (intersectional as well as additive) kinematic formulas [26, Chap. 5].
  • They can be additively extended to polyconvex sets (finite unions of compact, convex sets) and play a key role in the classical theory of geometric valuations (see [54, Chap. 1] and [26, Chap. 4.5]).
The intrinsic volumes have also been called Minkowski functionals (or quermassintegrals, though with a different normalization and notation). Due to their properties, the Minkowski functionals are useful descriptors that can be applied for the analysis of complex spatial structure.
Since the Minkowski functionals are isometry invariant, their usefulness in characterizing anisotropic features of objects under investigation is limited. The Minkowski tensors (see [25, Chap. 2]), which we introduce next, provide a more general set of tensor-valued geometric descriptors that are sensitive to position and orientation of geometric objects in Euclidean space. We start by providing a brief introduction to symmetric tensors and tensor-valued functionals as needed for the present purpose.
In the following, we use the scalar product of ${\mathbb{R}}^{d}$ to identify ${\mathbb{R}}^{d}$ with its dual space, hence a vector $a\in{\mathbb{R}}^{d}$ will be identified with the linear functional $x\mapsto\langle a,x\rangle$ from ${\mathbb{R}}^{d}$ to ${\mathbb{R}}$. For $r\in{\mathbb{N}}_{0}$, an $r$-tensor, or tensor of rank $r$, on ${\mathbb{R}}^{d}$ is an $r$-linear mapping from $({\mathbb{R}}^{d})^{r}$ to ${\mathbb{R}}$. The vector space of all $r$-tensors of ${\mathbb{R}}^{d}$ has dimension $d^{r}$. If $e_{1},\ldots,e_{d}$ is the standard basis of ${\mathbb{R}}^{d}$, then the $r$-tensors $e_{i_{1}}\otimes\cdots\otimes e_{i_{r}}$, $1\leq i_{1},\ldots,i_{r}\leq d$, are a basis of the vector space of all tensors of rank $r$ in ${\mathbb{R}}^{d}$. Here we have $e_{i_{1}}\otimes\cdots\otimes e_{i_{r}}(x_{1},\ldots,x_{r})=\prod_{j=1}^{r}\langle e_{i_{j}},x_{j}\rangle$ for $x_{1},\ldots,x_{r}\in{\mathbb{R}}^{d}$.
A tensor is symmetric if it is invariant under permutations of its arguments. By $\mathbb{T}^{r}({\mathbb{R}}^{d})$, or simply by $\mathbb{T}^{r}$, we denote the real vector space (with its standard topology) of symmetric $r$-tensors on ${\mathbb{R}}^{d}$. We define $\mathbb{T}^{0}={\mathbb{R}}$, and, by the identification made above, we have $\mathbb{T}^{1}={\mathbb{R}}^{d}$. In any case, $\dim\mathbb{T}^{r}({\mathbb{R}}^{d})=\binom{d+r-1}{r}$.
The symmetric tensor product $a_{1}\odot\dots\odot a_{k}\in\mathbb{T}^{r_{1}+\dots+r_{k}}$ of $a_{i}\in\mathbb{T}^{r_{i}}$, $i=1,\dots,k$, is a symmetric tensor of rank $r_{1}+\cdots+r_{k}$. Denoting by ${\mathcal{S}}(m)$ the group of bijections of $\{1,\ldots,m\}$, $s_{0}:=0$ and $s_{i}:=r_{1}+\dots+r_{i}$, for $i=1,\dots,k$, it is defined by
$(a_{1}\odot\dots\odot a_{k})(x_{1},\dots,x_{s_{k}}):=\frac{1}{s_{k}!}\sum_{\sigma\in{\mathcal{S}}(s_{k})}\prod_{i=1}^{k}a_{i}(x_{\sigma(s_{i-1}+1)},\dots,x_{\sigma(s_{i})})$
for $x_{1},\dots,x_{s_{k}}\in{\mathbb{R}}^{d}$. In this way, the space of symmetric tensors (of arbitrary rank) becomes an associative, commutative algebra with unit. For symmetric tensors $a,b,a_{i}$ we will use the abbreviations $a\odot b=:ab$ and $a_{1}\odot\dots\odot a_{k}=:a_{1}\cdots a_{k}$, and we write $a^{r}$ for the $r$-fold symmetric tensor product of $a$, if $r\in{\mathbb{N}}$, and $a^{0}:=1$. If $r_{1}=\ldots=r_{k}=1$ and $x_{1},\ldots,x_{k}\in{\mathbb{R}}^{d}$, then
$a_{1}\cdots a_{k}(x_{1},\ldots,x_{k})=\frac{1}{k!}\sum_{\sigma\in\mathcal{S}(k)}\prod_{i=1}^{k}\langle a_{i},x_{\sigma(i)}\rangle,$
and hence, for $a\in{\mathbb{R}}^{d}$ and $r\geq 1$, the $r$-fold symmetric tensor product of $a$ satisfies
$a^{r}(x_{1},\dots,x_{r})=\langle a,x_{1}\rangle\cdots\langle a,x_{r}\rangle.$
The scalar product, $Q(x,y)=\langle x,y\rangle$, $x,y\in{\mathbb{R}}^{d}$, is a symmetric tensor of rank two; we call $Q$ the metric tensor. Clearly, if $e_{1},\ldots,e_{d}$ is an orthonormal basis of ${\mathbb{R}}^{d}$, then $Q=e_{1}^{2}+\cdots+e_{d}^{2}$.
Let $(e_{1},\dots,e_{d})$ be an orthonormal basis of ${\mathbb{R}}^{d}$. Then the symmetric tensors $e_{i_{1}}\cdots e_{i_{r}}$ with $1\leq i_{1}\leq\dots\leq i_{r}\leq d$ form a basis of $\mathbb{T}^{r}({\mathbb{R}}^{d})$. The corresponding coordinate representation of $T\in\mathbb{T}^{r}({\mathbb{R}}^{d})$ is
$T=\sum_{1\leq i_{1}\leq\dots\leq i_{r}\leq d}t_{i_{1}\dots i_{r}}e_{i_{1}}\cdots e_{i_{r}}\quad\text{with}\quad t_{i_{1}\dots i_{r}}=\binom{r}{m_{1}\dots m_{d}}T(e_{i_{1}},\dots,e_{i_{r}}),$ (2.1)
where $m_{k}:=|\{\ell\in[r]:i_{\ell}=k\}|\in\{0,\ldots,r\}$, $k=1,\ldots,d$, counts how often the number $k$ appears among the indices $i_{1},\dots,i_{r}$. In particular, for $d=r=2$, we have $T(e_{1},e_{1})=t_{11}$, $T(e_{2},e_{2})=t_{22}$ and $T(e_{1},e_{2})=T(e_{2},e_{1})=\frac{1}{2}t_{12}$, hence $T=T(e_{1},e_{1})e_{1}^{2}+2T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{2})e_{2}^{2}$ with $2T(e_{1},e_{2})e_{1}e_{2}=T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{1})e_{2}e_{1}$.
For $T\in\mathbb{T}^{r}({\mathbb{R}}^{d})$ we introduce a norm by
$|T|:=\sup\{|T(x_{1},\ldots,x_{r})|:x_{i}\in{\mathbb{R}}^{d},\|x_{i}\|\leq 1,i=1,\ldots,r\}.$
If $a_{i}\in\mathbb{T}^{r_{i}}$, $i=1,\dots,k$, then
$|a_{1}\odot\dots\odot a_{k}|\leq|a_{1}|\cdots|a_{k}|.$ (2.2)
In particular, if $x\in{\mathbb{R}}^{d}$ and $u\in\mathbb{S}^{d-1}$, then $|x^{r}u^{s}|\leq\|x\|^{r}$. Moreover, if $x_{1},\ldots,x_{m}\in{\mathbb{R}}^{d}$ and $y_{1},\ldots,y_{m}\in{\mathbb{R}}^{d}$, then
$\left|\bigodot_{i=1}^{m}x_{i}-\bigodot_{i=1}^{m}y_{i}\right|\leq\sum_{j=1}^{m}\|x_{j}-y_{j}\|\prod_{i=1}^{j-1}\|x_{i}\|\prod_{i=j+1}^{m}\|y_{i}\|.$ (2.3)
For $T\in\mathbb{T}^{r}({\mathbb{R}}^{d})$ and a rotation $\vartheta\in O(d)$ of ${\mathbb{R}}^{d}$, we define the operation of $\vartheta$ on $T$ in such a way that the resulting tensor $\vartheta T\in\mathbb{T}^{r}({\mathbb{R}}^{d})$ is given by $(\vartheta T)(x_{1},\ldots,x_{r})=T(\vartheta^{-1}x_{1},\ldots,\vartheta^{-1}x_{r})$ for $x_{1},\ldots,x_{r}\in{\mathbb{R}}^{d}$.
In order to introduce the basic Minkowski tensors of (nonempty) convex bodies, we use the support measures which can be considered as local versions of the intrinsic volumes. For a convex body $K\subset{\mathbb{R}}^{d}$, the support measure $\Lambda_{j}(K,\cdot)$ is a Borel measure on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ which satisfies a local Steiner formula and $\Lambda_{j}(K,{\mathbb{R}}^{d}\times\mathbb{S}^{d-1})=V_{j}(K)$ for $j\in\{0,\ldots,d-1\}$. For basic properties of the support measures and a representation in the case where $K$ is a polytope, we refer to [53, Chapter 4.2]. The maps $K\mapsto\Lambda_{j}(K,\cdot)$ are measure-valued, additive (valuations), weakly continuous and for Borel sets $\alpha\subseteq{\mathbb{R}}^{d}$ and $\beta\subseteq\mathbb{S}^{d-1}$ they satisfy the covariance condition $\Lambda_{j}(gK,g\alpha\times g_{0}\beta)=\Lambda_{j}(K,\alpha\times\beta)$, where $g:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a rigid motion and $g_{0}$ is its rotational part (that is, $g(x)=g_{0}(x)+t$ for $x\in{\mathbb{R}}^{d}$, where $t\in{\mathbb{R}}^{d}$ is a translation vector and $g_{0}$ is a rotation). More generally, the support measures have been defined (as signed measures) for the more general class of sets of positive reach (see [49, Chapter 4] for an introduction) and for finite unions of compact sets with positive reach all of whose finite intersections have again positive reach (see [49, Chapter 5] and the literature cited there); the corresponding class of sets in ${\mathbb{R}}^{d}$ is denoted by $\mathcal{U}^{d}$. In fact, extensions of these measures to arbitrary closed subsets of ${\mathbb{R}}^{d}$ have been studied in [23, 24] and will be considered in Section 3.
If $K$ is a convex body or a compact set with positive reach, then the basic Minkowski tensors of $K$ are defined by
$\Phi_{k}^{r,s}(K):=\frac{1}{r!s!}\frac{\omega_{d-k}}{\omega_{d-k+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\Lambda_{k}(K,\mathrm{d}(x,u))$ (2.4)
for $k=0,\dots,d-1$ and $r,s\in{\mathbb{N}}_{0}$,
$\Phi_{d}^{r,0}(K):=\frac{1}{r!}\int_{K}x^{r}\,\mathrm{d}x,$ (2.5)
and by
$\Phi_{k}^{r,s}(K):=0\qquad\text{if $k\notin\{0,\dots,d\}$ or $r\notin{\mathbb{N}}_{0}$ or $s\notin{\mathbb{N}}_{0}$ or $k=d$, $s\not=0$}.$
Note that the tensor product $x^{r}u^{s}$ in the integrand is the symmetric tensor product of the symmetric tensor powers $x^{r}$ and $u^{s}$.
Minkowski tensors can be defined not only for convex bodies or compact sets with positive reach. By additivity their domain can be extended to the class of polyconvex sets (finite unions of convex bodies). Moreover they can be defined also for sets from the class $\mathcal{U}^{d}$, since this generalization is also true for the support measures. Furthermore, Minkowski tensors can be introduced for even more general classes of sets provided that the support measures are defined in a natural way and the integrals exist (see Section 3 and the references given there). An explicit description of the Minkowski tensors of polytopes is recalled in Section 4.4, in Section 4.5 we evaluate formulas for nonconvex sets. In the case of sets with positive reach or for even more general classes of compact sets $K$ (such as arbitrary finite unions of sets with positive reach), the support measures $\Lambda_{j}(K,\cdot)$ are consistently replaced by the reach measures $\mu_{j}(K;\cdot)$ in definition (2.4); the relationship between these measures will be discussed in Section 3.
It is a straightforward consequence of the properties of the support measures of convex bodies that the Minkowski tensors of convex bodies are continuous, isometry (rigid motion) covariant (for a detailed definition, especially of the underlying notion of polynomial behavior with respect to translations, we refer to [25, Section 2.2]) and additive tensor functionals (valuations) on the space of convex bodies in ${\mathbb{R}}^{d}$. By a fundamental result due to Alesker, the vector space $T_{p}$ of all mappings $\Gamma:\mathcal{K}^{d}\to\mathbb{T}^{p}$ having these properties is spanned by the tensor valuations $Q^{m}\Phi_{k}^{r,s}$, where $m,r,s\in{\mathbb{N}}_{0}$ satisfy $2m+r+s=p$, where $k\in\{0,\dots,d\}$, and where $s=0$ if $k=d$. In contrast to the real-valued and vector-valued case, these tensor valuations are no longer linearly independent. A study of linear dependencies was initiated by McMullen and completed by Hug, Schneider, Schuster (for precise references and detailed statements, see, e.g., [25, Thms. 2.6 and 2.7]). These general investigations imply that if $T_{2,k}$ denotes the subspace of all maps in $T_{2}$ that are homogeneous of degree $k$, then $T_{2}=T_{2,0}\oplus\cdots\oplus T_{2,d+2}$ (direct sum decomposition), $\dim T_{2}=3d+1$, and a basis of $T_{2,k}$ is displayed in Table 1. Denoting by $T_{2}^{\ast}$ the subspace of translation invariant maps in $T_{2}$ and by $T_{2,k}^{\ast}$ the subspace of translation invariant maps in $T_{2,k}$, we have $T_{2}^{\ast}=T_{2,0}^{\ast}\oplus\cdots\oplus T_{2,d}^{\ast}$, $\dim T_{2}^{\ast}=2d$, and a basis of $T_{2,k}^{\ast}$ is displayed in Table 1.
Table 1: Bases for the subspaces $T_{2,k}$, $k\in\{0,\ldots,d+2\}$, and $T_{2,k}^{\ast}$, $k\in\{0,\ldots,d\}$
$k=0$ $k=1$ $k\in\{2,\ldots,d-1\}$ $k=d$ $k\in\{d+1,d+2\}$
$T_{2,k}$$\{Q\Phi_{0}^{0,0}\}$$\{\Phi_{1}^{0,2},Q\Phi_{1}^{0,0}\}$$\{\Phi_{k}^{0,2},Q\Phi_{k}^{0,0},\Phi_{k-2}^{2,0}\}$$\{\Phi_{d-2}^{2,0},Q\Phi_{d}^{0,0}\}$$\{Q\Phi_{k-2}^{2,0}\}$
$T_{2,k}^{\ast}$$\{Q\Phi_{0}^{0,0}\}$$\{\Phi_{1}^{0,2},Q\Phi_{1}^{0,0}\}$$\{\Phi_{k}^{0,2},Q\Phi_{k}^{0,0}\}$$\{Q\Phi_{d}^{0,0}\}$
3 Asymptotics for estimators: a theoretical foundation
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. The topological boundary of $A$ is denoted by $\partial A$. The (positive) normal bundle of $A$ is the subset of ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ which is defined by
$\operatorname{Nor}(A):=\{(x,u)\in\partial A\times\mathbb{S}^{d-1}:{\rm dist}(A,x+ru)=r\text{ for some }r>0\}$
and the reach function of $A$ is given by
$r_{A}(x,u):=\sup\{s>0:{\rm dist}(A,x+su)=s\},\quad(x,u)\in\operatorname{Nor}(A)$
(note the slight correction in comparison with [23], see also [24]). In the following, we write $\mathcal{H}^{k}$ for the $k$-dimensional Hausdorff measure for $k\in\{0,\ldots,d\}$. We remark that $A$ has reach at least $r>0$ if $r_{A}(x,u)\geq r$ for $\mathcal{H}^{d-1}$-almost all $(x,u)\in\operatorname{Nor}(A)$; see [24, Lemma 3.19] for a more general fact.
It was shown in [23, Theorem 2.1], [24, Theorem 3.16] that there exist uniquely determined signed “reach measures” $\mu_{i}(A;\cdot)$ on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$, for $i\in\{0,\ldots,d-1\}$, concentrated on $\operatorname{Nor}(A)$ and such that
$\int_{{\mathbb{R}}^{d}\setminus A}f(z)\,\mathcal{H}^{d}(\mathrm{d}z)$
$\qquad=\sum_{i=0}^{d-1}\omega_{d-i}\int_{0}^{\infty}\int_{\operatorname{Nor}(A)}t^{d-1-i}\mathds{1}\{r_{A}(x,u)>t\}f(x+tu)\,\mu_{i}(A;\mathrm{d}(x,u))\,\mathrm{d}t,$ (3.1)
where $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is an arbitrary measurable bounded function with compact support. Clearly, (3) remains true if $f$ takes values in the (finite-dimensional) vector space $\mathbb{T}^{r}({\mathbb{R}}^{d})$, for any fixed $r\in{\mathbb{N}}_{0}$, where $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is said to be bounded if the map $x\mapsto|f(x)|$ is bounded ($f$ is bounded on a subset $B$ if the restriction of this map to $B$ is bounded).
Remark 3.1.
If $A\in\mathcal{U}^{d}$, then the support measures $\Lambda_{j}(A,\cdot)$ of $A$ are defined by additive extension. For this generalization it is crucial that $A$ can be obtained as a finite union of compact sets $A_{i}$ with positive reach, $i\in\{1,\ldots,m\}$, such that arbitrary finite intersections of these sets have again positive reach (compare [49, Example 4.16]). It was shown in [23, Section 3] that
$\mu_{j}(A;\cdot)=\Lambda_{j}(A,\operatorname{Nor}(A)\cap\cdot),\quad\text{for }j=0,\ldots,d-1.$
Moreover, the arguments in [23, Sections 3 and 4] show that indeed
$\mu_{d-1}(A;\cdot)=\Lambda_{d-1}(A,\cdot)$ (3.2)
as Borel measures on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$; see also [48, Corollary 2]. In particular, $A\mapsto\mu_{d-1}(A;\cdot)$ is additive on $\mathcal{U}^{d}$. On the other hand, the support measures are in general not defined on the larger domain of finite unions of sets with positive reach. For this reason, we work with the reach measures whenever the support measures are not available. Relation (3.2) shows that this approach is consistent with the previous literature if we restrict ourselves to the top order measures with index $j=d-1$ as in definition (4.2) below.
The integrals on the right-hand side of (3) are well defined, but in general the reach measures $\mu_{i}(A;\cdot)$ are not Radon measures. In the following, we consider compact sets $A\subset{\mathbb{R}}^{d}$ and assume that the (nonnegative) total variation measure $\|\mu_{i}(A)\|(\cdot)$ of $\mu_{i}(A;\cdot)$ is finite for $i\in\{0,\ldots,d-1\}$. A sufficient condition is provided in Lemma 3.2.
Lemma 3.2.
If $A$ is a finite union of compact sets of positive reach, then the total variation measures of the reach measures of $A$ are concentrated on $\operatorname{Nor}(A)$ and satisfy
$\|\mu_{i}(A)\|({\mathbb{R}}^{d}\times\mathbb{S}^{d-1})<\infty\quad\text{ for }i\in\{0,\ldots,d-1\}.$ (3.3)
Proof.
Recall that $\mu_{i}(A;\cdot)$ is concentrated on $\operatorname{Nor}(A)$ which implies the first assertion of the lemma.
It follows from [23, Corollary 2.5] that
$\|\mu_{i}(A)\|(\operatorname{Nor}(A))\leq c(d,i)\int_{\operatorname{Nor}(A)}|H_{d-1-i}(A,x,u)|\,\mathcal{H}^{d-1}(\mathrm{d}(x,u)),$
where $c(d,i)$ is a finite constant and $|H_{d-1-i}(A,x,u)|\leq\binom{d-1}{i}$, as can be seen from [23, (2.13)]. Let $A=\bigcup_{i=1}^{m}A_{i}$ with compact sets $A_{i}$ with positive reach. It is well known that $\operatorname{Nor}(A_{i})$ is bilipschitz homeomorphic to the compact submanifold ${\rm dist}(A_{i},\cdot)^{-1}(\{r_{0}\})$, if $r_{0}$ is smaller than the reach of $A_{i}$. Hence $\mathcal{H}^{d-1}(\operatorname{Nor}(A_{i}))<\infty$. Since
$\operatorname{Nor}(A)\subset\bigcup_{i=1}^{m}\operatorname{Nor}(A_{i}),$
the assertion follows. ∎
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. For $\mathcal{H}^{d}$-almost all $z\in{\mathbb{R}}^{d}\setminus A$, the metric projection $p_{A}(z)\in A$ to $A$ and the vector $u_{A}(z)\in\mathbb{S}^{d-1}$ are uniquely determined by
$\|p_{A}(z)-z\|={\rm dist}(A,z):=\min\{\|a-z\|:a\in A\},\quad u_{A}(z)=\frac{z-p_{A}(z)}{{\rm dist}(A,z)}.$
Note that $p_{A}(z)=z$ for $z\in A$, and if $(x,u)\in\operatorname{Nor}(A)$ and $0<s<r_{A}(x,u)$, then $p_{A}(x+su)=x$ and $u_{A}(x+su)=u$. If $A$ is a nonempty compact set, we consider
$\partial^{+}A:=\{x\in\partial A:(x,u)\in\operatorname{Nor}(A)\text{ for some }u\in\mathbb{S}^{d-1}\},$
and for $x\in\partial^{+}A$ we define the set of unit normal vectors of $A$ at $x$ by
$N(A,x):=\{u\in\mathbb{S}^{d-1}:(x,u)\in\operatorname{Nor}(A)\}.$
Note that $N(A,x)$ is either a singleton, consists of two unit vectors {±$\{\pm u\}$ or is an (infinite) spherically convex set. Then we set
$\partial^{i}A:=\{x\in\partial^{+}A:|N(A,x)|=i\},\quad i\in\{1,2\}.$
We write $\nu_{A}(x)$ for the unique unit vector in $N(A,x)$, for $x\in\partial^{1}A$, and ±$\pm\nu_{A}(x)$ for the two unit vectors in $N(A,x)$, if $x\in\partial^{2}A$. For $\varepsilon>0$, let $A^{\varepsilon}:=\{x\in{\mathbb{R}}^{d}:{\rm dist}(A,x)\leq\varepsilon\}$ denote the $\varepsilon$ neighborhood of $A$.
Theorem 3.3.
Let $r,s\in\mathbb{N}_{0}$. Let $A\subset{\mathbb{R}}^{d}$ be a nonemtpy compact set such that (3.3) holds. If $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is measurable and bounded on $\partial A$, then
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\int_{A^{\varepsilon}\setminus A}f(p_{A}(z))u_{A}(z)^{s}\,\mathcal{H}^{d}(\mathrm{d}z)$
$=\int_{\partial^{1}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x)+\left(1+(-1)^{s}\right)\int_{\partial^{2}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x).$
In particular, the assertion holds for finite unions of compact sets with positive reach.
Proof.
We apply (3) with the map $\tilde{f}:{\mathbb{R}}^{d}\to\mathbb{T}^{r+s}$, $\tilde{f}(z)=f(p_{A}(z))u_{A}(z)^{s}\mathds{1}\{z\in A^{\varepsilon}\setminus A\}$, which yields
$\int_{A^{\varepsilon}\setminus A}f(p_{A}(z))u_{A}(z)^{s}\,\mathcal{H}^{d}(\mathrm{d}z)$
$=\sum_{i=0}^{d-1}\omega_{d-i}\int_{0}^{\varepsilon}\int_{\operatorname{Nor}(A)}t^{d-1-i}f(x)u^{s}\mathds{1}\{r_{A}(x,u)>t\}\,\mu_{i}(A;\mathrm{d}(x,u))\,\mathrm{d}t.$
We consider
$h_{i}(t):=\omega_{d-i}t^{d-1-i}\int_{\operatorname{Nor}(A)}f(x)u^{s}\mathds{1}\{r_{A}(x,u)>t\}\,\mu_{i}(A;\mathrm{d}(x,u)).$
If $i\in\{0,\ldots,d-2\}$, then
$\varepsilon^{-1}\left|\int_{0}^{\varepsilon}h_{i}(t)\,\mathrm{d}t\right|$ $\leq\frac{\omega_{d-i}}{\varepsilon}\int_{0}^{\varepsilon}t^{d-1-i}\int_{\operatorname{Nor}(A)}|f(x)|\mathds{1}\{r_{A}(x,u)>t\}\,\|\mu_{i}(A)\|(\mathrm{d}{(}x,u))\,\mathrm{d}{t}$
$\leq\frac{\omega_{d-i}}{d-i}\cdot|f|_{\partial A}\cdot\|\mu_{i}(A)\|(\operatorname{Nor}(A))\cdot\varepsilon^{d-i-1}\to 0\quad\text{ as }\varepsilon\to 0_{+},$
where $|f|_{\partial A}:=\sup\{|f(x)|:x\in\partial A\}<\infty$ and basic properties of the norm ||$\cdot$|| ⋅ | on (symmetric) tensors were used, in particular (2.2). Moreover, by the dominated convergence theorem,
$h_{d-1}(t)$ $=2\int_{\operatorname{Nor}(A)}f(x)u^{s}\mathds{1}\{r_{A}(x,u)>t\}\,\mu_{d-1}(A;\mathrm{d}(x,u))$
$\to 2\int_{\operatorname{Nor}(A)}f(x)u^{s}\,\mu_{d-1}(A;\mathrm{d}(x,u))=:h_{d-1}(0_{+})\quad\text{ as }t\to 0_{+},$
since $r_{A}(x,u)>0$ for $(x,u)\in\operatorname{Nor}(A)$. Note that by [23, Proposition 4.1], $\mu_{d-1}(A;\cdot)$ is nonnegative and
$2\mu_{d-1}(A;\cdot)$ $=\int_{\partial^{1}A}\mathds{1}\{(x,\nu_{A}(x))\in\cdot\}\,\mathcal{H}^{d-1}(\mathrm{d}x)$ (3.4)
$\quad+\int_{\partial^{2}A}\mathds{1}\{(x,\nu_{A}(x))\in\cdot\}+\mathds{1}\{(x,-\nu_{A}(x))\in\cdot\}\,\mathcal{H}^{d-1}(\mathrm{d}x).$
Hence, if $t\to 0_{+}$, then
$h_{d-1}(t)$ $\to\int_{\partial^{1}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x)+\left(1+(-1)^{s}\right)\int_{\partial^{2}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x).$
Since
$\varepsilon^{-1}\int_{0}^{\varepsilon}h_{d-1}(t)\,\mathrm{d}t\to h_{d-1}(0_{+})\quad\text{ as }\varepsilon\to 0_{+},$
the assertion follows. ∎
By essentially the same argument, we obtain the following variant of Theorem 3.3. The case $s=0$ of the next theorem is already covered by Theorem 3.3 with the domain of integration restricted to the complement of the set $A$.
Theorem 3.4.
Let $s\in\mathbb{N}$. Let $A\subset{\mathbb{R}}^{d}$ be a nonemtpy compact set such that (3.3) holds. If $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is measurable and bounded on $\partial A$, then
$\lim_{\varepsilon\to 0_{+}}\frac{1+s}{\varepsilon^{1+s}}\int_{A^{\varepsilon}}f(p_{A}(z))(z-p_{A}(z))^{s}\,\mathcal{H}^{d}(\mathrm{d}z)$ (3.5)
$=\int_{\partial^{1}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x)+\left(1+(-1)^{s}\right)\int_{\partial^{2}A}f(x)\nu_{A}(x)^{s}\,\mathcal{H}^{d-1}(\mathrm{d}x).$
In particular, the assertion holds for finite unions of compact sets with positive reach.
Proof.
First, note that $z-p_{A}(z)=o$ if $z\in A$ and $s\neq 0$. Hence the domain of integration on the left-hand side can be restricted to $A^{\varepsilon}\setminus A$ without changing the value of the integral. We apply (3) with the map $\hat{f}:{\mathbb{R}}^{d}\to\mathbb{T}^{r+s}$, $\hat{f}(z)=f(p_{A}(z))(z-p_{A}(z))^{s}\mathds{1}\{z\in A^{\varepsilon}\setminus A\}$ and use that $(x+tu)-p_{A}(x+tu)=tu$ for $(x,u)\in\operatorname{Nor}(A)$ and $0<t<r_{A}(x,u)$. Thus we obtain
$\int_{A^{\varepsilon}}f(p_{A}(z))(z-p_{A}(z))^{s}\,\mathcal{H}^{d}(\mathrm{d}z)$
$=\sum_{i=0}^{d-1}\omega_{d-i}\int_{0}^{\varepsilon}\int_{\operatorname{Nor}(A)}t^{d-1-i+s}f(x)u^{s}\mathds{1}\{r_{A}(x,u)>t\}\,\mu_{i}(A;\mathrm{d}(x,u))\,\mathrm{d}t.$
Note that if $g:[0,\infty)\to{\mathbb{R}}$ is measurable and bounded, and $g(t)\to g(0_{+})$ as $t\to 0_{+}$, then
$\varepsilon^{-(1+s)}\int_{0}^{\varepsilon}t^{s}g(t)\,\mathrm{d}t\to\frac{g(0_{+})}{1+s}\quad\text{ as }\varepsilon\to 0_{+}.$
Otherwise, the proof follows the lines of the proof of Theorem 3.3. ∎
4 Algorithms to estimate Minkowski tensors
In this section, we define explicit algorithms that are used to estimate the Minkowski tensors, given a discrete (pixelated) data set. We focus on the current improvements in comparison to the original algorithm suggested in [21]. The digitized estimator of the volume tensors and especially the least squares fit with respect to the radial dependence strongly improve the robustness of the algorithm. Then we discuss details of our open-source implementation [47] and provide some advice on the choice of parameters. Finally, we demonstrate the performance for some test cases; see Sections 4.4 and 4.5.
4.1 Description and analysis
In the following, we describe two methods for estimating Minkowski tensors. The first is based on Corollary 4.1 and works for arbitrary finite unions of compact sets with positive reach and (surface) tensor functionals $\Phi^{r,s}_{d-1}$ with $s\geq 1$. The second approach, which is applicable for arbitrary tensor functionals $\Phi^{r,s}_{j}$, is based on solving a least squares problem and offers practical advantages. However, its theoretical foundation is only available for sets with positive reach. Nevertheless, in Section 4.5, we will see that it provides surprisingly accurate results for a set without positive reach.
To explain the two approaches, we fix integers $r,s\in{\mathbb{N}}_{0}$. The Voronoi tensor $\mathcal{V}_{R}^{r,s}(K)$ with distance parameter $R\geq 0$ and rank parameter tuple $(r,s)\in{\mathbb{N}}_{0}^{2}$ of a nonempty compact set $K\subset{\mathbb{R}}^{d}$ is defined by
$\mathcal{V}_{R}^{r,s}(K):=\int_{K^{R}}p_{K}(x)^{r}(x-p_{K}(x))^{s}\,\mathcal{H}^{d}(\mathrm{d}{x}),$ (4.1)
where $K^{R}$ is the set of points with distance at most $R>0$ from $K$ and $p_{K}$ is the (almost everywhere uniquely defined) metric projection on $K$. For $r=s=0$ the Voronoi tensor with distance parameter $R$ is the volume of $K^{R}$, i.e., the volume of the set of points with distance at most $R$ from $K$.
The following corollary of Theorem 3.4 shows how surface tensors can be approximated by Voronoi tensors. Recall the definition of the reach measure $\mu_{d-1}(A;\cdot)$ of a nonempty compact subset of ${\mathbb{R}}^{d}$ from Section 3, which equals the support measure $\Lambda_{d-1}(A,\cdot)$ if $A\in\mathcal{U}^{d}$ (see Remark 3.1 which ensures that the Minkowski tensor $\Phi^{r,s}_{d-1}$ in (4.2) is consistently defined for all $r,s\in{\mathbb{N}}_{0}$). An explicit description of the reach measure $\mu_{d-1}(A;\cdot)$ is provided in (3.4) (see [23, Proposition 4.1]).
Corollary 4.1.
Let $K\subset{\mathbb{R}}^{d}$ be a finite union of compact sets of positive reach. If $r\in{\mathbb{N}}_{0}$ and $s\in{\mathbb{N}}$, then
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon^{1+s}}\mathcal{V}^{r,s}_{\varepsilon}(K)=r!s!\kappa_{s+1}\Phi^{r,s}_{d-1}(K)$
and
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)\right)=r!2\,\Phi^{r,0}_{d-1}(K),$
where, for $r,s\in{\mathbb{N}}_{0}$,
$\Phi^{r,s}_{d-1}(K):=\frac{1}{r!s!}\frac{2}{\omega_{1+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}(x,u)).$ (4.2)
Proof.
First, we consider the case where $s\in{\mathbb{N}}$. We apply Theorem 3.4 with $f(x)=x^{r}$, $x\in{\mathbb{R}}^{d}$. Using the definition (4.1) of the total Voronoi tensor measure, we get the required expression on the left-hand side, after division by $1+s$. Accordingly, the right-hand side of equation (3.5) turns into
$\frac{2}{s+1}\int x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}{(}x,u)),$
where also relation (3.4) was used. The assertion now follows from the definition of the Minkowski tensor $\Phi^{r,s}_{d-1}(K)$.
If $s=0$, then an application of Theorem 3.3 and relation (3.4) yield
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)\right)$ $=\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{0}(K)\right)-\lim_{\varepsilon\to 0_{+}}\frac{\varepsilon}{\varepsilon^{2}}\left(\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)-\mathcal{V}^{r,0}_{0}(K)\right)$
$=2r!\Phi^{r,0}_{d-1}(K)-0,$
which proves the assertion. ∎
For a finite set $K_{0}\subset{\mathbb{R}}^{d}$ and $r,s\in{\mathbb{N}}_{0}$, the Voronoi tensors simplify to
$\mathcal{V}_{R}^{r,s}(K_{0})=\sum_{x\in K_{0}}x^{r}\int_{B(x,R)\cap V_{x}(K_{0})}(y-x)^{s}\,\mathcal{H}^{d}(\mathrm{d}{y}),$ (4.3)
where $B(x,R)$ is the closed ball with center $x$ and radius $R$ and
$V_{x}(K_{0}):=\{y\in{\mathbb{R}}^{d}:p_{K_{0}}(y)=x\}$
is the Voronoi cell of $x\in K_{0}$ with respect to the set $K_{0}$. It is clear that each compact set $K\subset{\mathbb{R}}^{d}$ can be approximated with any specified precision by finite sets $K_{0}\subset{\mathbb{R}}^{d}$ with respect to the Hausdorff metric $d_{H}$ on compact subsets of ${\mathbb{R}}^{d}$ (see, e.g., (4.10)). If $K,K_{0}\subseteq B(o,\rho)$ for some $\rho>0$, then
$\left|\mathcal{V}_{R}^{r,s}(K)-\mathcal{V}_{R}^{r,s}(K_{0})\right|\leq C(d,R,\rho,r,s)\cdot d_{H}(K,K_{0})^{\frac{1}{2}},$ (4.4)
which follows from [21, Thm. 4.1], where $C(d,R,\rho,r,s)$ is a constant depending only on the arguments in brackets. An inspection of the proof of Theorem 4.1 in [21] shows that the constant $C(d,R,\rho,r,s)$ can be chosen such that $C(d,R,\rho,r,s)\leq C(d,R_{0},\rho,r,s)$ if $0<R\leq R_{0}$. Note that (4.4) does not require that $K_{0}$ is a subset of $K$ (although an approximation of $K$ from inside can be guaranteed for a given $K$) and holds whenever $K,K_{0}$ are compact subsets of $B(o,\rho)$. Recall from [21] that $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
Evaluating the integral in (4.3) can be cumbersome and inefficient. Moreover, we do not require a precision beyond the approximation of $K$ via $K_{0}$. We, therefore, estimate the Voronoi tensors in (4.3) by considering a random grid
$\eta=\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\delta_{z+U_{a}}$ (4.5)
with some scaling parameter $a>0$ and a uniform random vector $U_{a}\sim\mathcal{U}([0,a]^{d})$. The tensor-valued random variable
$\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)=a^{d}\sum_{x\in\eta}\mathds{1}\big{\{}\|x-p_{K_{0}}(x)\|<R\big{\}}p_{K_{0}}(x)^{r}(x-p_{K_{0}}(x))^{s}$ (4.6)
will then be used as an estimator for ${\mathcal{V}}_{R}^{r,s}(K_{0})$. Since $K_{0}$ is a finite set, only finitely many summands do not vanish. The estimated Voronoi tensor in (4.6) can be viewed as a discrete version of the integral in (4.3). Figure 1 (a) provides an example data set $K_{0}$, which is given as a finite subset of an underlying set with positive reach. In Figure 1 (b) the idea of the estimator (4.6) is demonstrated for this example data set $K_{0}$. For each data point $x\in K_{0}$ one can see the Voronoi cell $V_{x}(K_{0})$.
Figure 1: Example data $K_{0}$ (red points), which is a finite subset of an underlying set with positive reach (a), and the Voronoi diagram (b) of the data $K_{0}$ and the set $K_{0}^{R}$ (blue) of points with distance not bigger than some $R>0$ from $K_{0}$. The algorithm uses a random grid $\eta$ (compare (4.6)) to estimate the Voronoi tensor $\mathcal{V}_{R}^{r,s}(K_{0})$ of $K_{0}$, where only the points of $\eta$ inside $K_{0}^{R}$ are relevant for the estimation.
Lemma 4.2.
Let $K_{0}\subset{\mathbb{R}}^{d}$ be a finite set and $a>0$. Then the estimator $\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ for ${\mathcal{V}}_{R}^{r,s}(K_{0})$ is unbiased, that is, $\mathbb{E}\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)={\mathcal{V}}_{R}^{r,s}(K_{0})$.
Proof.
Let $\{z_{i}:i\in{\mathbb{N}}\}=a\cdot{\mathbb{Z}}^{d}$ be an enumeration of the points in the lattice $a\cdot{\mathbb{Z}}^{d}$. Then
$\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ $=a^{d}\sum_{i\geq 1}\mathds{1}\{\|z_{i}+U_{a}-p_{K_{0}}(z_{i}+U_{a})\|<R\}p_{K_{0}}(z_{i}+U_{a})^{r}(z_{i}+U_{a}-p_{K_{0}}(z_{i}+U_{a}))^{s}$
$=a^{d}\sum_{i\geq 1}\sum_{w\in K_{0}}\mathds{1}\{p_{K_{0}}(z_{i}+U_{a})=w\}\mathds{1}\{\|z_{i}+U_{a}-w\|<R\}w^{r}(z_{i}+U_{a}-w)^{s}.$
Note that $p_{K_{0}}(z_{i}+U_{a})$ is uniquely defined $\mathbb{P}$-almost surely, since $U_{a}$ has a density with respect to the Lebesgue measure.
If the indicator functions are nonzero, then for every $i\in{\mathbb{N}}$ there is some $w\in K_{0}$ such that $\|z_{i}+U_{a}-w\|\leq R$, and hence $z_{i}\in K_{0}+B(\sqrt{d}a+R)$. In particular, this shows that the effective ranges of the summations over $i\geq 1$ and $w\in K_{0}$ are finite and can be chosen independently of $U_{a}$. Moreover, (2.2) yields
$|w^{r}(z_{i}+U_{a}-w)^{s}|\leq\|w\|^{r}\cdot\|z_{i}+U_{a}-w\|^{s}\leq C(K_{0})^{r}R^{s}<\infty,$
where $C(K_{0})$ is a constant that depends only on $K_{0}$, but not on a specific $w\in K_{0}$.
These remarks allow us to interchange summation and expectation so that
$\mathbb{E}\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ $=a^{d}\sum_{i\geq 1}\sum_{w\in K_{0}}\mathbb{E}\mathds{1}\{p_{K_{0}}(z_{i}+U_{a})=w\}\mathds{1}\{\|z_{i}+U_{a}-w\|<R\}w^{r}(z_{i}+U_{a}-w)^{s}$
$=\sum_{w\in K_{0}}\sum_{i\geq 1}\int_{z_{i}+[0,a]^{d}}\mathds{1}\{p_{K_{0}}(y)=w\}\mathds{1}\{\|y-w\|<R\}w^{r}(y-w)^{s}\,\mathcal{H}^{d}(\mathrm{d}y)$
$=\sum_{w\in K_{0}}\int_{{\mathbb{R}}^{d}}\mathds{1}\{y\in V_{w}(K_{0})\}\mathds{1}\{\|y-w\|<R\}w^{r}(y-w)^{s}\,\mathcal{H}^{d}(\mathrm{d}y)$
$=\sum_{w\in K_{0}}w^{r}\int_{{\mathbb{R}}^{d}}\mathds{1}\{y\in V_{w}(K_{0})\cap B(w,R)\}(y-w)^{s}\,\mathcal{H}^{d}(\mathrm{d}y)=\mathcal{V}_{R}^{r,s}(K_{0}),$
which proves the assertion. ∎
Next we show that $\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ is a consistent estimator for $\mathcal{V}_{R}^{r,s}(K)$, as $K_{0}\to K$ in the Hausdorff metric and $a=a(K_{0})\to 0$. More precisely, we obtain the following quantitative estimate. We write
$\operatorname{{\sf N}}(K_{0})=\min\{\|x_{1}-x_{2}\|:x_{1},x_{2}\in K_{0},x_{1}\neq x_{2}\}$
for the minimal distance of any two different points in a finite set $K_{0}\subset{\mathbb{R}}^{d}$ of cardinality at least two (which we tacitly assume all the time) and $s\wedge t$ for the minimum of $s,t\in{\mathbb{R}}$.
Theorem 4.3.
Let $K\subset{\mathbb{R}}^{d}$ be a compact set, let $r,s\in{\mathbb{N}}_{0}$ and $R>0$. Let $\rho>0$ be such that $K\subseteq B(o,\rho)$. If $K_{0}\subseteq B(o,\rho)$ is a finite set and $a>0$, then almost surely
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K)|\leq C^{\prime}(d,R,\rho,r,s)\left(d_{H}(K,K_{0})^{\frac{1}{2}}+\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}\right),$
where $C^{\prime}(d,R,\rho,r,s)$ is a positive constant, and $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
In particular, if $(K_{0}(i))_{i\in{\mathbb{N}}}$ is a sequence of finite sets in ${\mathbb{R}}^{d}$ and $(a_{i})_{i\in{\mathbb{N}}}$ is a sequence of positive numbers such that $K_{0}(i)\to K$ in the Hausdorff metric and $\operatorname{{\sf N}}(K_{0}(i))^{-1}a_{i}\to 0$, as $i\to\infty$, then $\widehat{\mathcal{V}}_{R}^{r,s}(K_{0}(i),a_{i})\to\mathcal{V}_{R}^{r,s}(K)$, as $i\to\infty$, almost surely.
Proof.
The constants employed in the proof are denoted by $c,c_{1},c_{2}$ and depend only the parameters indicated in brackets.
By the triangle inequality and in view of (4.4) it is sufficient to show that for a fixed finite set $K_{0}$, almost surely we have
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K_{0})|\leq c(d,R,\rho,r,s)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}.$
For this, observe that almost surely $p_{K_{0}}(z+U_{a})$ is uniquely defined for each $z\in a\cdot{\mathbb{Z}}^{d}$. Hence, almost surely
$\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ $=a^{d}\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}\{\|z+U_{a}-p_{K_{0}}(z+U_{a})\|<R\}p_{K_{0}}(z+U_{a})^{r}(z+U_{a}-p_{K_{0}}(z+U_{a}))^{s}$
$=\sum_{x\in K_{0}}a^{d}\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}\{\|z+U_{a}-x\|<R\}\mathds{1}\{z+U_{a}\in{V}_{x}(K_{0})\}x^{r}(z+U_{a}-x)^{s}$
$=\sum_{x\in K_{0}}x^{r}\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(z+U_{a})(z+U_{a}-x)^{s}a^{d}$
$=\sum_{x\in K_{0}}x^{r}\int\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(z+U_{a})(z+U_{a}-x)^{s}\mathds{1}_{z+[0,a)^{d}}(y)\,\mathcal{H}^{d}(\mathrm{d}y).$
Note that in the last step integration and summation can be interchanged, since for each $x\in K_{0}$ there are only finitely many summands $z\in a\cdot{\mathbb{Z}}^{d}$ such that the first indicator is nonzero. Moreover, we have
$\mathcal{V}_{R}^{r,s}(K_{0})$ $=\sum_{x\in K_{0}}x^{r}\int\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(y)(y-x)^{s}\,\mathcal{H}^{d}(\mathrm{d}y)$
$=\sum_{x\in K_{0}}x^{r}\int\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(y)(y-x)^{s}\mathds{1}_{z+[0,a)^{d}}(y)\,\mathcal{H}^{d}(\mathrm{d}y).$
Hence, using (2.2), we get
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K_{0})|$
$\leq\rho^{r}\sum_{x\in K_{0}}\int\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\left|\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(z+U_{a})(z+U_{a}-x)^{s}-\mathds{1}_{B(x,R)\cap{V}_{x}(K_{0})}(y)(y-x)^{s}\right|$
×$\qquad\qquad\qquad\qquad\times\mathds{1}_{z+[0,a)^{d}}(y)\,\mathcal{H}^{d}(\mathrm{d}y).$
If $z+[0,a)^{d}\subseteq B(x,R)\cap{V}_{x}(K_{0})$, then (2.2) and (2.3) shows that the expression ||$\cdot$|| ⋅ | under the integral can be bounded from above by
$\left|(z+U_{a}-x)^{s}-(y-x)^{s}\right|\leq s\cdot\|z+U_{a}-y\|\cdot R^{s-1}\leq s\cdot\sqrt{d}R^{s-1}\cdot a.$
If $(z+[0,a)^{d})\cap B(x,R)\cap{V}_{x}(K_{0})=\emptyset$, then the expression is zero. In the remaining case, we have $(z+[0,a)^{d})\cap\partial(B(x,R)\cap{V}_{x}(K_{0}))\neq\emptyset$ and the expression ||$\cdot$|| ⋅ | is bounded from above by $2R^{s}$ (see (2.2)). Thus we obtain
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K_{0})|$
$\leq\rho^{r}\sum_{x\in K_{0}}\left(\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}\{z+[0,a)^{d}\subseteq B(x,R)\cap{V}_{x}(K_{0})\}\cdot s\cdot\sqrt{d}R^{s-1}\cdot a\cdot a^{d}\right.$
$\qquad\qquad\qquad\qquad\left.+\sum_{z\in a\cdot{\mathbb{Z}}^{d}}\mathds{1}\{(z+[0,a)^{d})\cap\partial(B(x,R)\cap{V}_{x}(K_{0}))\neq\emptyset\}\cdot 2R^{s}\cdot a^{d}\right)\allowdisplaybreaks$
$\leq d\rho^{r}R^{s-1}\sum_{x\in K_{0}}\Big{(}s\cdot a\,\mathcal{H}^{d}(B(x,R)\cap{V}_{x}(K_{0}))$
$\qquad\qquad\qquad\qquad+R\,\mathcal{H}^{d}\left(\partial(B(x,R)\cap{V}_{x}(K_{0}))+B(o,\sqrt{d}\cdot a)\right)\Big{)}\allowdisplaybreaks$
$\leq s\cdot d\rho^{r}R^{s-1}\cdot a\cdot\mathcal{H}^{d}\left(K+B(o,d_{H}(K,K_{0})+R)\right)$
$\qquad\qquad\qquad\qquad+d\rho^{r}R^{s}\sum_{x\in K_{0}}\mathcal{H}^{d}\left(\partial(B(x,R)\cap{V}_{x}(K_{0}))+B(o,\sqrt{d}\cdot a)\right),$ (4.7)
where we used that the Voronoi cells have pairwise disjoint interiors. Since $K,K_{0}\subset B(o,\rho)$, we get
$\mathcal{H}^{d}(K+B(o,d_{H}(K,K_{0})+R))\leq\mathcal{H}^{d}\left(B(o,2\rho+R)\right)\leq\kappa_{d}(2\rho+R)^{d}.$ (4.8)
To deal with the remaining sum, we denote by $r_{x}$ the inradius of the Voronoi cell $V_{x}(K_{0})$. Using a volume bound from the proof of Lemma 3.6 in [22] and the obvious inequality $r_{x}\geq\operatorname{{\sf N}}(K_{0})/2$ for $x\in K_{0}$, we obtain
$\mathcal{H}^{d}\left(\partial(B(x,R)\cap{V}_{x}(K_{0}))+B(o,\sqrt{d}\cdot a)\right)$
$\leq 2\left[\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})+B(o,\sqrt{d}\cdot a)\right)-\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)\right]$
$\leq 2\left[\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})+\frac{\sqrt{d}\cdot a}{r_{x}\wedge R}\cdot(B(x,R)\cap{V}_{x}(K_{0}))\right)-\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)\right]$
$=\left[\left(1+\frac{\sqrt{d}\cdot a}{r_{x}\wedge R}\right)^{d}-1\right]\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$
$\leq c_{1}(d)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})\wedge R},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})\wedge R}\right)^{d}\right\}\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right),$
and hence
$\mathcal{H}^{d}\left(\partial(B(x,R)\cap{V}_{x}(K_{0}))+B(o,\sqrt{d}\cdot a)\right)$
$\leq c_{2}(d,\rho,R)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right),$ (4.9)
where we use that $\operatorname{{\sf N}}(K_{0})\leq 2\rho$ if $R<\operatorname{{\sf N}}(K_{0})$. If $R\geq\operatorname{{\sf N}}(K_{0})$, then we can choose $c_{2}(d,\rho,R)=c_{1}(d)$. The sum of the volumes $\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$ over $x\in K_{0}$ is bounded from above by $\kappa_{d}(\rho+R)^{d}$. Combination of (4.1), (4.8), and (4.1) finally yields
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K_{0})|$
$\leq d\rho^{r}R^{s-1}(2\rho+R)^{d}\cdot\left(s\cdot a+R\cdot c_{2}(d,\rho,R)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}\right),$
which implies the required bound, since $\operatorname{{\sf N}}(K_{0})\leq 2\rho$. ∎
From the comments after (4.4) and the proof of Theorem 4.3, which yields more explicit information about the constants involved, we get the following consequence.
Corollary 4.4.
Let $K\subset{\mathbb{R}}^{d}$ be a compact set, and let $r,s\in{\mathbb{N}}_{0}$. Let $\rho>0$ be such that $K\subseteq B(o,\rho)$. If $K_{0}\subseteq B(o,\rho)$ is a finite set, $R\leq R_{0}$, and $0<a\leq\operatorname{{\sf N}}(K_{0})\wedge R$, then almost surely
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K)|\leq C^{\prime\prime}(d,R_{0},\rho,r,s)\left(d_{H}(K,K_{0})^{\frac{1}{2}}+\frac{a}{\operatorname{{\sf N}}(K_{0})\wedge R}\right),$
where $C^{\prime\prime}(d,R_{0},\rho,r,s)>0$ is a constant depending only on $d,R_{0},\rho,r,s$, and $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
Combination of Corollary 4.1 and Corollary 4.4 finally yields the next result. It serves as a theoretical foundation for the approximation of surface tensors of finite unions of sets with positive reach via Corollary 4.1. We remark that for any nonempty compact set $K\subset{\mathbb{R}}^{d}$, the finite set
$K(n):=\left\{(\sqrt{d}n)^{-1}\cdot z:z\in{\mathbb{Z}}^{d},{\rm dist}(K,(\sqrt{d}n)^{-1}\cdot z)\leq\frac{1}{n}\right\}$ (4.10)
satisfies $d_{H}(K(n),K)\leq\frac{1}{n}$.
Theorem 4.5.
Let $K\subset{\mathbb{R}}^{d}$ be a finite union of compact sets with positive reach, and let $r\in{\mathbb{N}}_{0}$, $s\in{\mathbb{N}}$. Let $\rho>0$ be such that $K\subseteq B(o,\rho)$. If $(\varepsilon_{n})_{n\in{\mathbb{N}}}$ satisfies $\varepsilon_{n}\to 0_{+}$ as $n\to\infty$, then there exist sequences $(K_{0}(n))_{n\in{\mathbb{N}}}$ of finite subsets of ${\mathbb{R}}^{d}$ with $K_{0}(n)\to K$ as $n\to\infty$ and $(a_{n})_{n\in{\mathbb{N}}}$ of real numbers such that, almost surely,
$\lim_{n\to\infty}\frac{\widehat{\mathcal{V}}_{\varepsilon_{n}}^{r,s}(K_{0}(n),a_{n})}{\varepsilon_{n}^{1+s}}=r!s!\kappa_{s+1}\Phi^{r,s}_{d-1}(K)$
and
$\lim_{n\to\infty}\frac{1}{\varepsilon_{n}}\left(\widehat{\mathcal{V}}_{\varepsilon_{n}}^{r,0}(K_{0}(n),a_{n})-\widehat{\mathcal{V}}_{\varepsilon_{n}^{2}}^{r,0}(K_{0}(n),a_{n})\right)=r!2\,\Phi^{r,0}_{d-1}(K).$
Proof.
We consider only the case $s=0$, the argument for $s\in{\mathbb{N}}$ is similar (but easier). By Corollary 4.1, for $n\in{\mathbb{N}}$ there is some $\varepsilon(n)>0$ such that
$\left|\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)\right)-r!2\,\Phi^{r,0}_{d-1}(K)\right|\leq\frac{1}{2n}\quad\text{if }0<\varepsilon\leq\varepsilon(n).$ (4.11)
where $\varepsilon(n+1)<\varepsilon(n)\leq\varepsilon(1)\leq 1$ for $n\in{\mathbb{N}}$ and $\varepsilon(n)\to 0_{+}$ as $n\to\infty$.
We can assume that $\varepsilon_{m}\leq\varepsilon(1)$ for $m\in{\mathbb{N}}$. For $m\in{\mathbb{N}}$ there is a unique $n\in{\mathbb{N}}$ such that $\varepsilon_{m}\in(\varepsilon(n+1),\varepsilon(n)]$. By the remarks preceding Theorem 4.5 , there is a finite set $K_{0}(m)\subset{\mathbb{R}}^{d}$ such that
$d_{H}(K,K_{0}(m))\leq\left(C^{\prime\prime}(d,\varepsilon(1),\rho,r,0)8n\right)^{-2}\cdot\varepsilon_{m}^{2}.$ (4.12)
Moreover, we choose
$a_{m}\leq\frac{\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}^{2}}{C^{\prime\prime}(d,\varepsilon(1),\rho,r,0)8n}\cdot\varepsilon_{m},$ (4.13)
where we can assume that $C^{\prime\prime}(d,\varepsilon(1),\rho,r,0)4n\geq 1$, so that
$0<a_{m}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}^{2}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}\leq 1$
Corollary 4.4, (4.12) and (4.13) then yield almost surely
$\left|\widehat{\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K_{0}(m),a_{m})-\mathcal{V}^{r,0}_{\varepsilon_{m}}(K)\right|\leq\frac{\varepsilon_{m}}{4n},\quad\left|\widehat{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K_{0}(m),a_{m})-\mathcal{V}^{r,0}_{\varepsilon_{m}^{2}}(K)\right|\leq\frac{\varepsilon_{m}}{4n},$
hence
$\left|\frac{1}{\varepsilon_{m}}\left(\widehat{\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K_{0}(m),a_{m})-\widehat{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K_{0}(m),a_{m})\right)-\frac{1}{\varepsilon_{m}}\left({\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K)-{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K)\right)\right|\leq\frac{1}{2n}.$ (4.14)
Combining (4.11) with $\varepsilon=\varepsilon_{m}$ and (4.14), we arrive at
$\left|\frac{1}{\varepsilon_{m}}\left(\widehat{\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K_{0}(m),a_{m})-\widehat{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K_{0}(m),a_{m})\right)-r!2\,\Phi^{r,0}_{d-1}(K)\right|\leq\frac{1}{n},$
which implies the assertion. ∎
Using the estimator from (4.6), Corollary 4.1 can in principle be applied directly to estimate the tensors $\Phi^{r,s}_{d-1}(K)$ from Corollary 4.1 by choosing a very small value for $R$, whenever $K$ is a finite union of compact sets with positive reach and $s\geq 1$. We tested the performance for a rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ (see Example 4.9), a spherical shell with inner radius 1 and outer radius 2 (see Example 4.10), and for a rectangle from which a smaller (open) rectangle has been removed (see Example 4.11). The generated finite test data are obtained by intersecting these objects with a grid. The simulated estimates can be found in Table 2. To achieve good results, however, we needed an $a$ that is significantly smaller than $R$, which itself must already be small. Since a small value of $a$ considerably slows down the computation of the estimator from (4.6), this is a substantial disadvantage in practice. Therefore, for estimating the tensors from Corollary 4.1 we propose to use the alternative approach via a least squares problem, whenever it is applicable. For example, a single rendition (i.e., a run of the algorithm that provides an estimate) for the spherical shell $S$ in the simulations for Table 2 took about 21 minutes on a standard personal computer, whereas in the simulations from Table 5 with the approach via a least squares problem, all 25 renditions for the same spherical shell took only about 3 minutes on a standard personal computer. Note that two renditions differ in the grid (4.5) used, as randomness only comes into play here. Surprisingly, the alternative approach led to excellent results even in cases where we only have a heuristic justification (see Table 6 for Example 4.11). Finally, we point out that we observed that the second part of Theorem 4.5, namely the approach for $s=0$, requires far too much runtime and memory capacity for practical applications. Since $a$ must be significantly smaller than $\varepsilon$ and $\varepsilon^{2}$ for the application of the Theorem, and $\varepsilon$ itself should already be a very small value, $a$ becomes too small (and thus the number of points in the lattice process too large) for practical use. However, for $r=s=0$ this problem can be overcome, as the following remark demonstrates.
Remark 4.6.
Let $K$ be a finite union of compact sets with positive reach and $\operatorname{tr}(T)$ the trace of a tensor $T\in\mathbb{T}^{2}({\mathbb{R}}^{d})$, that is, $\operatorname{tr}(T):=\sum_{j=1}^{d}T(e_{j},e_{j})$. Then it is true that
$\operatorname{tr}\big{(}\Phi_{d-1}^{0,2}(K)\big{)}=\frac{1}{4\pi}\Phi_{d-1}^{0,0}(K),$
where we used that $\operatorname{tr}(u^{2})=1$ for $u\in\mathbb{S}^{d-1}$ and the linearity of the trace operator.
Table 2: Results for the 2-dimensional rectangle $K=[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$, the spherical shell $S$ with inner radius 1 and outer radius 2 (see Example 4.10) and $R_{a,b}$ from Example 4.11 with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$ by using the estimator from (4.6) and the result from Corollary 4.1. We intersected $K,S,R_{a,b}$ with a grid of resolution $0.0005$ and choose $R=\varepsilon=0.05$ and $a=0.0005$. We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Estimates Tensors Values Estimates
[…]
($\Phi_{1}^{0,2}$ ($K$))2,2 0.2387 0.2438 ($\Phi_{1}^{0,2}$ ($S$))1,1 0.3754 0.3754
[…]
($\Phi_{1}^{0,2}$ ($R_{a,b}$))2,2 0.3173 0.3178
In the following, we explain the algorithm that we will focus on for the remainder of the paper. The algorithm is based on the idea from [21] (see Section 3.2 there) to estimate the Minkowski tensors of a set $K$ with positive reach by first approximating $K$ by a simpler set $K_{0}$ (for instance, $K_{0}$ could be a finite subset of $K$) and by solving a linear system $Ax=b$. The solution $x$ of the system is the $(d+1)$-dimensional vector containing the estimated Minkowski tensors $\hat{\Phi}_{d}^{r,s}(K_{0}),\dots,\hat{\Phi}_{0}^{r,s}(K_{0})$, where $r,s\in{\mathbb{N}}_{0}$. The vector $b$ contains the estimated Voronoi tensors $\hat{\mathcal{V}}_{R_{0}}^{r,s}(K_{0}),\dots,\hat{\mathcal{V}}_{R_{d}}^{r,s}(K_{0})$, for $d+1$ different values $0<R_{0}<\dots<R_{d}$ of the radius, and the matrix $A$ is a Vandermonde-type matrix. It is shown in [21] that the solution of this system yields the exact values of the Minkowski tensors whenever the exact values of the Voronoi tensors are plugged in and $R_{d}$ is smaller than the reach of $K$. Since inverting a matrix is numerically not stable, we replace the matrix inversion by a least squares problem, which allows us to use more than only $d+1$ equations. However, the basic idea of our algorithm remains to estimate the Minkowski tensors of $K$ via the Voronoi tensors of a finite set $K_{0}$. Here $K_{0}$ can be viewed as a finite approximation of a possibly infinite set $K$. In practice $K_{0}$ is usually considered to be a finite sample of points in $K$, while for generating test data one can intersect $K$ with a grid.
Now fix $a>0$, an integer $n\geq d+1$, and $0<R_{1}<\dots<R_{n}$. One of the key ideas of the algorithm is to replace the matrix inversion from [21] by a linear least squares problem. Extending the linear system from [21, (6)], we define the estimators $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$, $i\in\{0,\dots,d\}$, as the solution of the linear least squares problem
$\widehat{\mathcal{V}}_{R_{i}}^{r,s}(K_{0},a)=\sum_{k=0}^{d}r!s!\kappa_{k+s}R_{i}^{k+s}\widehat{\Phi}_{d-k}^{r,s}(K_{0},a,R),\qquad i\in\{1,\ldots,n\},$ (4.15)
where $R:=(R_{1},\dots,R_{n})$. Recall that $\kappa_{n}$ denotes the volume of an $n$-dimensional unit ball. Moreover, a symmetric tensor $T\in\mathbb{T}^{r+s}({\mathbb{R}}^{d})$ of rank $r+s$ is determined by the values $T(e_{i_{1}},\dots,e_{i_{r+s}})$ for $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ (see (2.1)). The algorithm computes the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$, $i\in\{0,\dots,d\}$, by solving the least squares problem from (4.15) for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$, using standard methods. The following lemma is well known (see [9, Theorems 1.1.5 and 1.1.6] or [20, Theorem 9]).
Lemma 4.7.
Let $s\in{\mathbb{N}}_{0}$ and $n\geq d+1$. Let $x_{1},y_{1},\dots,x_{n},y_{n}\in{\mathbb{R}}$, $c_{0},\dots,c_{d}\in{\mathbb{R}}\setminus\{0\}$ and
$X:=\left(c_{j}x_{i}^{s+j}\right)_{\begin{subarray}{c}i=1,\ldots,n\\ j=0,\ldots,d\end{subarray}}\in{\mathbb{R}}^{n\times(d+1)},\qquad y:=(y_{1},\ldots,y_{n})^{\top}\in{\mathbb{R}}^{n}.$ (4.16)
If $x_{i}\neq x_{j}$ for $i\neq j$, then $X$ has rank $d+1$ and $\mathbf{a}=(X^{\top}X)^{-1}X^{\top}y\in{\mathbb{R}}^{d+1}$ is the unique vector satisfying
$\mathbf{a}=\operatorname*{arg\,min}\Big{\{}\sum_{i=1}^{n}(b_{0}c_{0}x_{i}^{s}+\dots+b_{d}c_{d}x_{i}^{s+d}-y_{i})^{2}\mid(b_{0},\dots,b_{d})^{\top}\in{\mathbb{R}}^{d+1}\Big{\}}.$
Remark 4.8.
If $s\geq 1$, then we can also consider the matrix in ${\mathbb{R}}^{n\times d}$ with $n\geq d$ obtained from $X$ by deleting the first column and setting $b_{0}=0$, which corresponds to the definition $\Phi_{d}^{r,s}=0$ if $s\neq 0$.
As a consequence of Lemma 4.7 the least squares problem from (4.15) is continuous, that is whenever the input vector converges to some vector $b\in{\mathbb{R}}^{n}$, the solution will converge to the solution of the least squares problem with input $b$. In particular, if $K$ is a set with positive reach and $R_{n}$ is less then the reach of $K$, then the tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$ from (4.15) converge to ${\Phi}_{j}^{r,s}(K)$ as $\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)$ converges to $\mathcal{V}_{R}^{r,s}(K)$, for which Theorem 4.3 provides sufficient conditions. Note that the approximating tensors $\widehat{\Phi}_{j}^{r,s}(K_{0},a,R)$ depend on the choice of the fixed radii $0<R_{1}<\cdots<R_{n}$, but the Minkowski tensors ${\Phi}_{j}^{r,s}(K)$ will be independent of the radii (as they should be) as long as $K$ has reach greater than $R_{n}$. We would like to point out that although the solution to the least squares problem from Lemma 4.7 involves a matrix inversion, the algorithm for solving the problem does not compute an inverse matrix. Lemma 4.7 merely serves to show that the solution is continuous with respect to the input. In our situation we apply Lemma 4.7 for each choice of $1\leq i_{1}\leq\dots\leq i_{r+s}\leq d$ to
$x_{i}=R_{i},\quad c_{i}=r!s!\kappa_{i+s},\quad y_{i}=\widehat{\mathcal{V}}_{R_{i}}^{r,s}(K_{0},a)(e_{i_{1}},\dots,e_{i_{r+s}}),$
to obtain $\widehat{\Phi}_{d-j}^{r,s}(K_{0},a,R)(e_{i_{1}},\dots,e_{i_{r+s}})$ for $j\in\{0,\dots,d\}$. So there are two main tasks involved in the algorithm: To estimate the Voronoi tensors of $K_{0}$ with respect to the $n$ values $R_{1},\dots,R_{n}$ by the estimator in (4.6) and then to solve the system (4.15). Therefore the only randomness involved in the algorithm is contained in the random grid $\eta$.
We now explain the role of the relevant quantities. As an input the algorithm takes a finite set $K_{0}$, the rank parameters $r,s\in{\mathbb{N}}_{0}$ of the Minkowski tensors, the number $n\geq d+1$ of equations as well as the maximal radius $R_{n}$. The smallest radius $R_{1}$ used is the average nearest neighbor distance in $K_{0}$, that is
$av(K_{0}):=\frac{1}{|K_{0}|}\sum_{x\in K_{0}}\min\{\|x-z\|:z\in K_{0}\setminus\{x\}\},$
where $|K_{0}|$ denotes the cardinality of $K_{0}$. The other radii are chosen equidistantly between $R_{1}$ and $R_{n}$. Therefore we have the additional condition that $R_{n}$ has to be greater than $av(K_{0})$, and we recommend to choose $R_{n}$ to be at least $(d+1)\cdot av(K_{0})$. The resolution $a$ of the random grid $\eta$ is also chosen as $av(K_{0})$.
It is also possible to provide the algorithm with an observation window
$W=\prod_{i=1}^{d}[a_{i},b_{i}],\quad a_{i}<b_{i}\text{ for }i\in\{1,\ldots,d\},$
instead of $R_{n}$. In this case the algorithm computes the distance of the data from the boundary of the observation window in each direction and and chooses the minimum of these values for $R_{n}$. By the distance of $K_{0}$ from the boundary of $W$ in some direction $s\cdot e_{j}$, where $s\in\{-1,1\}$ and $e_{j}$ denotes the $j$-th standard unit vector, we mean
$\min\{b_{j}-x_{j}:x=(x_{1},\dots,x_{d})^{\top}\in K_{0}\},\quad s=1,$
$\min\{x_{j}-a_{j}:x=(x_{1},\dots,x_{d})^{\top}\in K_{0}\},\quad s=-1.$
The output of the algorithm consists of a list containing the $d+1$ estimated Minkowski tensors $\widehat{\Phi}_{d}^{r,s}(K_{0},a,R),\dots,\widehat{\Phi}_{0}^{r,s}(K_{0},a,R)$ (in this order), where a tensor is given as a dictionary whose keys are the indices of the individual values. Note that the tensors are symmetric and therefore the algorithm does not estimate all the values since some of them are redundant. For a tuple of indices $(i_{1},\dots,i_{r+s})\in\{1,\ldots,d\}^{r+s}$ with $i_{1}\leq\dots\leq i_{r+s}$ the algorithm estimates the value of the tensor corresponding to these indices, but not those corresponding to permutations of the tuple since they are the same. Table 3 shows a small summary of the input of the algorithm.
Table 3: Input parameters of the algorithm
$K_{0}$ finite input data
$r,s$ rank parameters of the Minkowski tensors
$n$ number of equations considered ($n\geq d+1$)
$R_{n}$ maximal radius considered (we recommend $R_{n}\geq(d+1)R_{1}$)
$W$ observation window containing $K_{0}$ (optional instead of $R_{n}$)
4.2 Implementation
The algorithm is implemented in Python, and the code is available via an open-source package [47]. One can apply the algorithm by using the Python functions Voromink and Vorosurf in the file VorominkEstimation.py. Details can be found in the description of the code, which can also be found in [47]. Closely following the above described estimation of the Minkowski tensors, the user chooses some rank parameters $r,s\in{\mathbb{N}}_{0}$. Then the algorithm will return estimators of the $d+1$ Minkowski tensors $\Phi_{0}^{r,s}$$(K),\dots,$$\Phi_{d}^{r,s}$$(K)$, based on the input data $K_{0}$ and the dimension $d$, for a (in general) unknown set $K$. The algorithm also offers the option to rotate the random grid $\eta$ on which the algorithm is based. This option is useful for test data that lies somehow parallel to the axis spanned by the standard vectors. For real data this rotation should, in general, not be necessary and can be skipped to further reduce the run time (especially in high dimensions).
We found the run times of our code to be convenient for practically relevant examples. For example, a single rendition of the simulations for a two-dimensional rectangle (results can be found in Appendix A), containing about 150,000 points, took an average of 30 seconds. In this case, the algorithm was applied for 5 different rank parameter tuples $(r,s)$ (all possible combinations with $r+s\leq 2$), and for each tuple, 50 runs were performed (resulting in a total of 250 runs and a computation time of 1.5 hours). Computation time scales with the number of points of the grid process $\eta$. Therefore the time increases highly with the dimension $d$ and with a refinement of the resolution of the grid process which is taken as the average nearest neighbor distance $av(K_{0})$ of the input data $K_{0}$. According to Theorem 4.3, it would theoretically be better to choose the grid resolution even smaller than $av(K_{0})$. However, in our simulations, we found that this did not significantly improve the results but considerably increased the runtime.
4.3 Choice of parameters
The choices of the parameters $n,R_{n}$ has a big influence on the results of the algorithm. Therefore we share the experience we made while working with the algorithm. For the parameter $n$ we obtained better results for higher values. In principle one can choose $n$ as big as possible, but the distances $R_{i+1}-R_{i}$ between the values of consecutive radii should not be smaller than the average nearest neighbor distance in the test data. We often used the value $n=50$.
The choice of the parameter $R_{n}$ is much more intricate. The theory requires the value of $R_{n}$ to be smaller than the reach of the original set $K$ that is represented by the data $K_{0}$. In practice this value is usually unknown. If the original set is supposed to be convex, $R_{n}$ could in principle be arbitrarily big. In the general case one should choose this parameter depending on an assessment of the data. It is important that $R_{n}$ is not too small. We recommend it to be at least $d+1$ times the average nearest neighbor distance $av(K_{0})$ in the test data $K_{0}$, but from our experiences this value sometimes needs to be increased further. For more advanced applications and requirements, Bayesian optimization can be a helpful tool for obtaining the best possible parameters.
4.4 Convex test cases
In this section we present some simulation results showing how the algorithm works for specific simulated test data. We tested the algorithm for rectangles in dimension $d\in\{2,3\}$ as well as for a rectangle in dimension 2 with rounded vertices. To generate the data we fixed a grid $a\cdot{\mathbb{Z}}^{d}$ and intersected it with the rectangle. Table 4 shows formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$. These formulas can be obtained by the following simplification of the Minkowski tensors of polytopes. For any polytope $P\subset{\mathbb{R}}^{d}$ we have
$\Phi_{k}^{r,s}(P)$ $=\frac{1}{r!s!}\frac{1}{\omega_{d-k+s}}\sum_{F\in\mathcal{F}_{k}(P)}\int_{F}\int_{N(P,F)\cap\mathbb{S}^{d-1}}x^{r}u^{s}\,\mathcal{H}^{d-k-1}(\mathrm{d}u)\,\mathcal{H}^{k}(\mathrm{d}x),\qquad k<d,$
$\Phi_{d}^{r,0}(P)$ $=\frac{1}{r!}\int_{P}x^{r}\,\mathrm{d}x,\qquad\qquad\qquad\Phi_{d}^{r,s}(P)=0,\qquad s>0,$
where $\mathcal{F}_{k}(P)$ is the set of the $k$-dimensional faces of the polytope $P$ and $N(P,F)$ is the normal cone of $P$ with respect to the face $F$. Recall that $\mathcal{H}^{k}$ denotes the $k$-dimensional Hausdorff measure and $\omega_{k}$ is the surface area of the unit sphere $\mathbb{S}^{k-1}$ in ${\mathbb{R}}^{k}$.
Example 4.9.
Let $R=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $a_{1}$ and $a_{2}$. Then the above formula yields for $r=2,s=0,k=1$ that
$\Phi_{1}^{2,0}(R)=\frac{1}{4}\sum_{F\in\mathcal{F}_{1}(R)}\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x).$
Using $x^{2}=x_{1}^{2}e_{1}^{2}+2x_{1}x_{2}e_{1}e_{2}+x_{2}^{2}e_{2}^{2}$, for an edge of the form $F=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times\{\pm\frac{a_{2}}{2}\}$ we obtain
$\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x)$ $=\int_{-\frac{a_{1}}{2}}^{\frac{a_{1}}{2}}y^{2}e_{1}^{2}\pm ya_{2}e_{1}e_{2}+\frac{a_{2}^{2}}{4}e_{2}^{2}\,\mathrm{d}y=\frac{a_{1}^{3}}{12}e_{1}^{2}+\frac{a_{1}a_{2}^{2}}{4}e_{2}^{2}.$
Therefore the summation over all the edges of $R$ yields
$\Phi_{1}^{2,0}(R)$ $=\frac{1}{24}\left((a_{1}^{3}+3a_{1}^{2}a_{2})e_{1}^{2}+(3a_{1}a_{2}^{2}+a_{2}^{3})e_{2}^{2}\right).$
For $r=s=k=1$ we obtain
$\Phi_{1}^{1,1}(R)$ $=\frac{1}{2\pi}\sum_{F\in\mathcal{F}_{1}(R)}\int_{F}\int_{N(R,F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x).$
As before we start by computing the integral for an edge of the form $F=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times\{\pm\frac{a_{2}}{2}\}$. The normal cone of this set intersected with the unit sphere is given by {±$\{\pm e_{2}\}$. Since the tensor $xe_{2}$ can be written as $x_{1}e_{1}e_{2}+x_{2}e_{2}^{2}$, we obtain
$\int_{F}\int_{N(R,F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x)$ =±$=\pm\int_{F}xe_{2}\,\mathcal{H}^{1}(\mathrm{d}x)=\pm\int_{-\frac{a_{1}}{2}}^{\frac{a_{1}}{2}}ye_{1}e_{2}\pm\frac{a_{2}}{2}e_{2}^{2}\,\mathrm{d}y=\frac{a_{1}a_{2}}{2}e_{2}^{2}.$
After taking the summation over all edges, we get
$\Phi_{1}^{1,1}(R)$ $=\frac{a_{1}a_{2}}{2\pi}(e_{1}^{2}+e_{2}^{2})=\frac{a_{1}a_{2}}{2\pi}\,Q.$
The results of the algorithm for a two-dimensional rectangle can be found in Appendix A and are very close to the exact values. We also tested a 3-dimensional rectangle in the same way and obtained very good results. Further we considered a 2-dimensional rectangle with “rounded corners” and the results again were very close to the exact values (relative deviations of less than 1% for rectangles of varying sizes with up to 1.3 million points). The detailed results, along with those for a 3-dimensional rectangle, where we once again achieved very good outcomes, can be found in Appendix A.
Table 4: Formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$, where the values of $\Phi_{k}^{1,0}$, $\Phi_{k}^{0,1}$ vanish for all $k\in\{0,1,2\}$.
Tensor Formula Tensor Formula Tensor Formula
$\Phi_{0}^{0,0}$ 1 $\Phi_{1}^{0,0}$ $a_{1}+a_{2}$ $\Phi_{2}^{0,0}$ $a_{1}\cdot a_{2}$
$\Phi_{0}^{0,2}$ $\frac{1}{4\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{1}^{0,2}$ $\frac{1}{4\pi}\left(a_{2}e_{1}^{2}+a_{1}e_{2}^{2}\right)$ $\Phi_{2}^{0,2}$ 0
$\Phi_{0}^{2,0}$ $\frac{1}{8}\left(a_{1}^{2}e_{1}^{2}+a_{2}^{2}e_{2}^{2}\right)$ $\Phi_{1}^{2,0}$ $\frac{1}{24}\sum_{i=1}^{2}\left(a_{i}^{3}+3a_{i}^{2}a_{j}\right)e_{i}^{2}$ $\Phi_{2}^{2,0}$ $\frac{1}{24}\left(a_{1}^{3}a_{2}e_{1}^{2}+a_{2}^{3}a_{1}e_{2}^{2}\right)$
$\Phi_{0}^{1,1}$ $\frac{1}{2\pi}\left(a_{1}e_{1}^{2}+a_{2}e_{2}^{2}\right)$ $\Phi_{1}^{1,1}$ $\frac{a_{1}a_{2}}{2\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{2}^{1,1}$ 0
4.5 Nonconvex test cases
If $A\subset{\mathbb{R}}^{d}$ is a set of positive reach, the boundary $\partial A$ of $A$ is of class $C^{2}$, and $A$ has a unique exterior unit outer normal vector $\nu(A,x)$ at each $x\in\partial A$, then
$\Lambda_{m}(A,\cdot)=\frac{1}{\omega_{d-m}}\int_{\partial A}\mathds{1}\{(x,\nu(A,x))\in\cdot\}\sum_{|I|=d-1-m}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}),$
where $k_{1}(A,x),\ldots,k_{d-1}(A,x)\in{\mathbb{R}}$ are the principal curvatures of $A$ at $x\in\partial A$ and the summation extends over all subsets $I\subset\{1,\ldots,d-1\}$ of cardinality $|I|=d-1-m$ (if $m=d-1$ the empty product is interpreted as $1$). Thus we obtain, for $k\in\{0,\dots,d-1\}$,
$\Phi^{r,s}_{k}(A)=\frac{1}{r!s!\omega_{d-k+s}}\int_{\partial A}x^{r}\nu(A,x)^{s}\sum_{|I|=d-1-m}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}).$
Example 4.10.
Let $A=\{x\in{\mathbb{R}}^{d}:\rho_{1}\leq\|x\|\leq\rho_{2}\}$ denote a spherical shell with inner radius $\rho_{1}$ and outer radius $\rho_{2}$, where $0\leq\rho_{1}<\rho_{2}<\infty$. Observe that the reach of $A$ is $\rho_{1}$ whenever $\rho_{1}>0$ (if $\rho_{1}=0$, then $A$ is a Euclidean ball with radius $\rho_{2}$). Since the principal curvatures of $A$ at $x\in\partial A$ are equal to $\rho_{2}^{-1}$, if $\|x\|=\rho_{2}$, and equal to $-\rho_{1}^{-1}$, if $\|x\|=\rho_{1}>0$, we simply write $k(A,x)$ for these principal curvatures and get
$\Phi^{r,s}_{k}(A)$ $=\frac{\binom{d-1}{k}}{r!s!\omega_{d-k+s}}\int_{\partial A}x^{r}\nu(A,x)^{s}k(A,x)^{d-1-k}\,\mathcal{H}^{d-1}(\mathrm{d}{x})$
$=\frac{\binom{d-1}{k}}{r!s!\omega_{d-k+s}}\int_{{\mathbb{S}}^{d-1}}\left((\rho_{2}u)^{r}u^{s}\rho_{2}^{k+1-d}\rho^{d-1}_{2}+(\rho_{1}u)^{r}(-u)^{s}(-\rho_{1})^{k+1-d}\rho^{d-1}_{1}\right)\,\mathcal{H}^{d-1}(\mathrm{d}{u})$
$=\frac{\binom{d-1}{k}}{r!s!\omega_{d-k+s}}\left(\rho_{2}^{r+k}+(-1)^{s+d-1-k}\rho_{1}^{r+k}\right)\int_{{\mathbb{S}}^{d-1}}u^{r+s}\,\mathcal{H}^{d-1}(\mathrm{d}{u})$
$=\mathds{1}\{r+s\text{ even}\}\frac{\binom{d-1}{k}2\omega_{d+r+s}}{r!s!\omega_{d-k+s}\omega_{r+s+1}}\left(\rho_{2}^{r+k}+(-1)^{s+d-1-k}\rho_{1}^{r+k}\right)Q^{\frac{r+s}{2}},$
where the transformation formula and [27, Lemma 7] have been used to determine the integral.
If $k=d-1$, this specializes to
$\Phi^{r,s}_{d-1}(A)=\mathds{1}\{r+s\text{ even}\}\frac{2\omega_{d+r+s}}{r!s!\omega_{s+1}\omega_{r+s+1}}\left(\rho_{2}^{r+d-1}+(-1)^{s}\rho_{1}^{r+d-1}\right)Q^{\frac{r+s}{2}},$
and if $k=d-1$ and $r=0$, we obtain
$\Phi^{0,s}_{d-1}(A)=\mathds{1}\{s\text{ even}\}\frac{2\omega_{d+s}}{s!\omega_{s+1}^{2}}\left(\rho_{1}^{d-1}+\rho_{2}^{d-1}\right)Q^{\frac{s}{2}}.$
Table 5 shows simulation results for the 2-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$.
Table 5: Results for the 2-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$ intersected with a grid of resolution $a=0.005$. The parameter choices were $n=50$ and $R_{n}=1.8$ (0.9 times the reach of the spherical shell). We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Algorithm Tensors Values Algorithm
[…]
$\Phi_{2}^{0,0}$ 9.4248 9.4018 ($\Phi_{1}^{0,2}$)1,1 0.3749 0.3749
[…]
($\Phi_{1}^{1,1}$)2,2 1.5 1.492
Example 4.11.
Let $R_{a}=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $a_{1}$ and $a_{2}$ and $a=(a_{1},a_{2})^{\top}\in{\mathbb{R}}^{2}$. For $0<a_{1}<b_{1}$ and $0<a_{2}<b_{2}$ we consider $R_{a,b}:=R_{b}\setminus R_{a}^{\circ}$, i.e., the interior of the rectangle $R_{a}$ is removed from the enclosing rectangle $R_{b}$. The set $R_{a,b}$ does not have positive reach, but it is a polyconvex set and in particular a finite union of compact sets with positive reach. The surface tensors $\Phi_{1}^{r,s}(R_{a,b})$ can be expressed in terms of the surfaces tensors of $R_{a}$ and $R_{b}$ by
$\Phi_{1}^{r,s}(R_{a,b})=\Phi_{1}^{r,s}(R_{b})+(-1)^{s}\Phi_{1}^{r,s}(R_{a}),$
where $r,s\in{\mathbb{N}}_{0}$; see Table 4 for explicit formulas. Table 6 shows simulation results for $R_{a,b}$ with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$.
Table 6: Results for $R_{a,b}$ with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=50$ and $R_{n}=0.45$ ($0.9\cdot 0.5$ times the length of the smaller side of $R_{a}$). We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Algorithm Tensors Values Algorithm
 $\Phi_{0}^{0,0}$ 0 -0.2837 $\Phi_{1}^{0,0}$ 11 11.01
$\Phi_{2}^{0,0}$ 13 13.00 ($\Phi_{1}^{0,2}$)1,1 0.5570 0.5477
[…]
($\Phi_{1}^{1,1}$)2,2 2.069 2.073
5 Isotropic random polytopes
Next, we apply our algorithm to a common model of random convex polytopes. In Section 5.1 we first show how the expected value of the Minkowski tensors $\Phi^{0,s}_{k}(Z)$ for an isotropic random polytope $Z$ can be expressed in terms of the metric tensor $Q$ and the expected value of the intrinsic volume $V_{k}(Z)$. From this relation we then derive explicit formulas for beta-polytopes, a class of isotropic random polytopes that has proved to be useful in stochastic geometry (see, e.g., [32]). In Section 5.2, the exact evaluations of the resulting formulas are compared to the results of a simulation study based on our algorithm.
5.1 Exact expectations
In this section, we focus on the case $r=0$. A random polytope $Z$ in ${\mathbb{R}}^{d}$ is said to be isotropic if $\rho Z$ and $Z$ have the same distribution for all $\rho\in SO(d)$.
Lemma 5.1.
Let $Z$ be an isotropic random polytope in ${\mathbb{R}}^{d}$. Let $k\in\{0,\ldots,d-1\}$. If $\mathbb{E}V_{k}(Z)<\infty$, then
$\mathbb{E}\Phi^{0,s}_{k}(Z)=\mathds{1}\{s\text{ \rm even}\}\frac{2\omega_{d+s}\omega_{d-k}}{s!\omega_{d}\omega_{s+1}\omega_{d-k+s}}\mathbb{E}V_{k}(Z)\cdot Q^{\frac{s}{2}};$
in particular, the surface tensors are given by
$\mathbb{E}\Phi^{0,s}_{d-1}(Z)=\mathds{1}\{s\text{ \rm even}\}\frac{4\omega_{d+s}}{s!\omega_{d}\omega_{s+1}^{2}}\mathbb{E}V_{d-1}(Z)\cdot Q^{\frac{s}{2}},$ (5.1)
Proof.
Let $\rho\in SO(d)$. Since $Z$ is isotropic, we deduce from the definition of $\Phi^{0,s}_{k}(Z)$ that
$\mathbb{E}\Phi^{0,s}_{k}(Z)$ $=\frac{1}{s!\omega_{d-k+s}}\mathbb{E}\sum_{F\in\mathcal{F}_{k}(\rho Z)}V_{k}(F)\int_{N(\rho Z,F)\cap{\mathbb{S}}^{d-1}}u^{s}\,\mathcal{H}^{d-k-1}(\mathrm{d}u)$
$=\frac{1}{s!\omega_{d-k+s}}\mathbb{E}\sum_{F\in\mathcal{F}_{k}(Z)}V_{k}(\rho F)\int_{\rho N(Z,F)\cap{\mathbb{S}}^{d-1}}u^{s}\,\mathcal{H}^{d-k-1}(\mathrm{d}u)$
$=\frac{1}{s!\omega_{d-k+s}}\mathbb{E}\sum_{F\in\mathcal{F}_{k}(Z)}V_{k}(F)\int_{N(Z,F)\cap{\mathbb{S}}^{d-1}}(\rho v)^{s}\,\mathcal{H}^{d-k-1}(\mathrm{d}v).$ (5.2)
Integration of (5.1) over all $\rho\in SO(d)$ with respect to the Haar probability measure $\nu$ on $SO(d)$ and Fubini’s theorem yield
$\mathbb{E}\Phi^{0,s}_{k}(Z)$ $=\frac{1}{s!\omega_{d-k+s}}\mathbb{E}\sum_{F\in\mathcal{F}_{k}(Z)}V_{k}(F)\int_{N(Z,F)\cap{\mathbb{S}}^{d-1}}\int_{SO(d)}(\rho v)^{s}\,\nu(\mathrm{d}\rho)\,\mathcal{H}^{d-k-1}(\mathrm{d}v).$
The application of Fubini’s theorem (in particular, interchanging expectation and integration) is justified since $\mathbb{E}V_{k}(Z)<\infty$. From
$\int_{SO(d)}(\rho v)^{s}\,\nu(\mathrm{d}\rho)=\frac{1}{\omega_{d}}\int_{{\mathbb{S}}^{d-1}}w^{s}\,\mathcal{H}^{d-1}(\mathrm{d}w)=\mathds{1}\{s\text{ \rm even}\}\frac{2\omega_{d+s}}{\omega_{d}\omega_{s+1}}Q^{\frac{s}{2}}$
(see, e.g., [27, Lemma 7] and the reference given there) we deduce that
$\mathbb{E}\Phi^{0,s}_{k}(Z)$ $=\mathds{1}\{s\text{ \rm even}\}\frac{2\omega_{d+s}}{s!\omega_{d}\omega_{s+1}\omega_{d-k+s}}Q^{\frac{s}{2}}\mathbb{E}\sum_{F\in\mathcal{F}_{k}(Z)}V_{k}(F)\mathcal{H}^{d-k-1}(N(Z,F)\cap{\mathbb{S}}^{d-1})$
$=\mathds{1}\{s\text{ \rm even}\}\frac{2\omega_{d+s}\omega_{d-k}}{s!\omega_{d}\omega_{s+1}\omega_{d-k+s}}\mathbb{E}V_{k}(Z)\cdot Q^{\frac{s}{2}},$
which proves the assertion. ∎
We consider the case where $Z=P^{\beta}_{\ell,d}$ is a beta-polytope with $\beta>-1$. To define this class of random polytopes, let
$f_{d,\beta}(x):=c_{d,\beta}\mathds{1}\{\|x\|<1\}\left(1-\left\|x\right\|^{2}\right)^{\beta},\quad x\in{\mathbb{R}}^{d},$
with the normalizing constant
$c_{d,\beta}:=\frac{\Gamma\left(\frac{d}{2}+\beta+1\right)}{\pi^{\frac{d}{2}}\Gamma\left(\beta+1\right)}.$
Hence $f_{d,\beta}$ is the density of a beta-distribution in ${\mathbb{R}}^{d}$.
Let $X_{1},\ldots,X_{\ell}$ be i.i.d. random points in ${\mathbb{R}}^{d}$ whose common distribution function has Lebesgue density $f_{d,\beta}$. If $\ell\geq d+1$, then
$P^{\beta}_{\ell,d}:=\operatorname{conv}\{X_{1},\ldots,X_{\ell}\}$
is called a beta-polytope (see [32]) in ${\mathbb{R}}^{d}$ (with parameters $\beta$ and $\ell$). Since the density $f_{d,\beta}$ is rotation invariant, the distribution of $X_{1},\ldots,X_{\ell}$ is rotation invariant, and hence $P^{\beta}_{\ell,d}$ is an isotropic random polytope. It follows from [31, Cor. 2.4] that
$\mathbb{E}V_{d-1}(P^{\beta}_{\ell,d})$ $=\frac{d(2\beta+d+1)}{2^{d}\Gamma\left(\frac{d}{2}\right)}\binom{\ell}{d}\left(\frac{\Gamma\left(\beta+\frac{d+2}{2}\right)}{\Gamma\left(\beta+\frac{d+3}{2}\right)}\right)^{d}\!\!\int_{-1}^{1}(1-h^{2})^{d\beta+\frac{(d-1)(d+2)}{2}}F_{1,\beta+\frac{d-1}{2}}(h)^{\ell-d}\,\mathrm{d}h$
with
$F_{1,\beta+\frac{d-1}{2}}(h):=\frac{\Gamma\left(\beta+\frac{d+2}{2}\right)}{\sqrt{\pi}\Gamma\left(\beta+\frac{d+1}{2}\right)}\int_{-1}^{h}(1-x^{2})^{\beta+\frac{d-1}{2}}\,\mathrm{d}x,\quad h\in[-1,1].$
This expectation can be rewritten in the form
$\mathbb{E}V_{d-1}(P^{\beta}_{\ell,d})$ $=\frac{d\pi^{\frac{d}{2}}}{(2\beta+d+1)^{d-1}\Gamma\left(\frac{d}{2}\right)}\binom{\ell}{d}\left(\frac{\Gamma\left(\beta+\frac{d+2}{2}\right)}{\sqrt{\pi}\Gamma\left(\beta+\frac{d+1}{2}\right)}\right)^{\ell}$
×$\qquad\times\int_{-1}^{1}(1-h^{2})^{d\beta+\frac{d-1}{2}(d+2)}\left(\int_{-1}^{h}(1-x^{2})^{\beta+\frac{d-1}{2}}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
If $\beta=-\frac{1}{2}$, then
$\mathbb{E}V_{d-1}(P^{-\frac{1}{2}}_{\ell,d})$ $=\frac{\pi^{\frac{d}{2}}}{d^{d-2}\Gamma\left(\frac{d}{2}\right)}\binom{\ell}{d}\left(\frac{\Gamma\left(\frac{d+1}{2}\right)}{\sqrt{\pi}\Gamma\left(\frac{d}{2}\right)}\right)^{\ell}\int_{-1}^{1}\sqrt{1-h^{2}}^{d^{2}-2}\left(\int_{-1}^{h}\sqrt{1-x^{2}}^{d-2}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
For $d=2$ and $\ell\geq 3$ this simplifies to
$\mathbb{E}V_{1}(P^{-\frac{1}{2}}_{\ell,2})=\pi\frac{\ell-1}{\ell+1}.$
If $\beta=\frac{1}{2}$, then
$\mathbb{E}V_{d-1}(P^{\frac{1}{2}}_{\ell,d})$ $=\frac{d\pi^{\frac{d}{2}}}{(d+2)^{d-1}\Gamma\left(\frac{d}{2}\right)}\binom{\ell}{d}\left(\frac{\Gamma\left(\frac{d+3}{2}\right)}{\sqrt{\pi}\Gamma\left(\frac{d+2}{2}\right)}\right)^{\ell}$
×$\qquad\times\int_{-1}^{1}\sqrt{1-h^{2}}^{d^{2}+2d-2}\left(\int_{-1}^{h}\sqrt{1-x^{2}}^{d}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
If $d=2$, this simplifies to
$\mathbb{E}V_{1}(P^{\frac{1}{2}}_{\ell,2})=\pi\cdot\ell(\ell-1)\frac{9}{4^{\ell+1}}\int_{-1}^{1}(1-h)^{3}(1+h)^{2\ell-1}(2-h)^{\ell-2}\,\mathrm{d}h.$
Expansion of $(2-h)^{\ell-2}=(1+(1-h))^{\ell-2}$ and repeated partial integration yield
$\mathbb{E}V_{1}(P^{\frac{1}{2}}_{\ell,2})=\pi\cdot 9(\ell-1)\sum_{j=0}^{\ell-2}\frac{\binom{\ell-2}{j}}{\binom{2\ell+j+3}{j+3}}2^{j}.$
From (5.1) we have
$\mathbb{E}\Phi^{0,s}_{d-1}(P^{\beta}_{\ell,d})=\mathds{1}\{s\text{ \rm even}\}\frac{4\omega_{d+s}}{s!\omega_{d}\omega_{s+1}^{2}}\mathbb{E}V_{d-1}(P^{\beta}_{\ell,d})\cdot Q^{\frac{s}{2}},$
with $\mathbb{E}V_{d-1}(P^{\beta}_{\ell,d})$ given as above.
Similar formulas can be derived for $\mathbb{E}\Phi^{0,s}_{k}(P^{\beta}_{\ell,d})$, where $k\in\{0,\ldots,d-1\}$, $s\in{\mathbb{N}}_{0}$ or $k=d$, $s=0$, from results provided in [31] in combination with Lemma 5.1. The same is true for some other classes of isotropic random polytopes, including beta-polytopes or symmetric versions of beta- and beta-polytopes. For example for the expected intrinsic volumes of a beta-polytope Proposition 2.3 in [31] yields
$\mathbb{E}V_{k}\left(P^{\beta}_{\ell,d}\right)=\binom{d}{k}\frac{\kappa_{d}}{\kappa_{k}\kappa_{d-k}}\mathbb{E}V_{k}\left(P^{\beta+\frac{d-k}{2}}_{\ell,k}\right).$
5.2 Simulation study
Table 7: Results for 2- and 4-dimensional beta-polytopes $P_{10,d}^{\beta}$. We intersected the simulated polytope with a grid of resolution 0.005 (resp. 0.02 for $d=4$). The parameter choices were $n=50$ and $R_{n}=1$ (resp. $R_{n}=0.5$ for $d=4$). We took the average of 50 (resp. $10$ for $d=4$) renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of the expectation $\mathbb{E}\left[V_{k}\left(P_{{10},{d}}^{\beta}\right)\right]$.
$d$ $k$ $\beta$ Exact values Algorithm $d$ $k$ $\beta$ Exact values Algorithm
$2$ $1$ $\frac{1}{2}$ 2.08… 2.05(3) $4$ $3$ $\frac{1}{2}$ 0.81… 0.84(4)
$2$ $1$ $-\frac{1}{2}$ 2.57… 2.57(3) $4$ $3$ $-\frac{1}{2}$ 1.35… 1.56(8)
$2$ $2$ $\frac{1}{2}$ 1.06… 1.03(3) $4$ $4$ $\frac{1}{2}$ 0.107… 0.093(8)
$2$ $2$ $-\frac{1}{2}$ 1.71… 1.72(6) $4$ $4$ $-\frac{1}{2}$ 0.21… 0.23(2)
In this section we simulate specific beta-polytopes in dimension 2 and 4 and determine the Minkowski tensors of the resulting polytopes from the implementation of our algorithm. Since the formulas from Section 5.1 simplify for $d=2$ and $\beta\in\{-\frac{1}{2},\frac{1}{2}\}$, we focus on these two values for $\beta$ (also for $d=4$). Table 7 shows the results for the expected volume and surface area of a 2- and a 4-dimensional beta-polytope with the choice $\ell=10$. For $d=2$ we also simulated the case $\ell=100$ and obtained similar results. Furthermore, Table 8 contains the results for the expectation of the tensor $\Phi_{1}^{0,2}(P_{10,2}^{\beta})_{i,j}$ (left) resp. $\Phi_{1}^{0,4}(P_{10,2}^{\beta})_{i,j,k,l}$, and Table 9 gives the results for the expectation of the tensor $\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}$.
Table 8: Results for 2-dimensional beta-polytopes $P_{10,2}^{\beta}$. We intersected the simulated polytope with a grid of resolution 0.005. The parameter choices were $n=50$ and $R_{n}=1$. We took the average of 50 (left) resp. 10 (right) renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of
$\mathbb{E}\left[\Phi_{1}^{0,2}(P_{10,2}^{\beta})_{i,j}\right]$ $\mathbb{E}\left[\Phi_{1}^{0,4}(P_{10,2}^{\beta})_{l,l,l,l}\right]$
$\beta$ $i$ $j$ Exact values Algorithm $\beta$ $l$ Exact values Algorithm
$\frac{1}{2}$ 1 1 0.083… 0.084(2) $\frac{1}{2}$ 1 0.0025… 0.0025(1)
$\frac{1}{2}$ 1 2 0 -0.002(2) $\frac{1}{2}$ 2 0.0025… 0.0023(1)
$\frac{1}{2}$ 2 2 0.083… 0.084(2) $-\frac{1}{2}$ 1 0.0031… 0.0030(1)
$-\frac{1}{2}$ 1 1 0.102… 0.102(2) $-\frac{1}{2}$ 2 0.0031… 0.0031(1)
$-\frac{1}{2}$ 1 2 0 0.001(2)
$-\frac{1}{2}$ 2 2 0.102… 0.101(2)
[…]
Table 9: Results for 4-dimensional beta-polytopes $P_{10,4}^{\beta}$. We intersected the simulated polytopes with a grid of resolution 0.02. The parameter choices were $n=50$ and $R_{n}=0.5$. We took the average of 10 renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of the expectation $\mathbb{E}\left[\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}\right]$.
$\beta$ $i$ $j$ Exact values Algorithm $\beta$ $i$ $j$ Exact values Algorithm
$\frac{1}{2}$ 1 1 0.016… 0.010(2) $-\frac{1}{2}$ 1 1 0.027… 0.022(3)
$\frac{1}{2}$ 2 2 0.016… 0.011(1) $-\frac{1}{2}$ 2 2 0.027… 0.020(3)
$\frac{1}{2}$ 3 3 0.016… 0.011(1) $-\frac{1}{2}$ 3 3 0.027… 0.022(4)
$\frac{1}{2}$ 4 4 0.016… 0.014(2) $-\frac{1}{2}$ 4 4 0.027… 0.019(2)
[…]
6 Experimental data
Finally, we apply our methods to real data, more specifically, to two distinct use cases. The first example is a collection of metallic grains that can be considered to be (or at least well approximated by) sets of positive reach. The second example are nanorough surfaces that exhibit deep cuts and high ridges, i.e., they are clearly no sets of positive reach, but they can be thought of as finite unions of compact sets with positive reach. By our Theorem 3.4, we can use our methods to estimate the interfacial tensor $\Phi_{1}^{0,2}$.
6.1 Metallic grains
Figure 2: Examples of three distinct metallic grains that are represented by voxel centers. Here, each voxel center is indicated by a ball with a diameter equal to the lattice spacing, i.e., 1 $\mu$m. The underlying experimental data of the grains is from [60].
Figure 3: Non-normalized histograms of the estimated volumes (a) and surface areas (b) of the metallic grains from Section 6.1. Besides a large number of small cells, we observe a few exceptionally large cells. For better visualization, only a range of values is shown, i.e., the most extreme cases are excluded here. The underlying experimental data of the grains is from [60].
Figure 4: Non-normalized histograms of anisotropy indices for the metallic grains from Section 6.1. More specifically, the plots show histograms of the ratios between the smallest and largest absolute eigenvalues of the interfacial tensor $\Phi_{1}^{0,2}$ (a) and the curvature tensor $\Phi_{2}^{0,2}$ (b). The underlying experimental data of the grains is from [60].
We first apply our algorithm and code to a real data set published in [60] of a polycrystalline material, more specifically, of a nickel-based superalloy; see [60], for further details. The data set set contains, among others, 2546 individual grains, each represented by a set of voxel centers; see Figure 2, for three examples. The resolution of the grid is 1 $\mu$m. To exclude numerical artifacts, we apply a data cut and only consider grains with at least 500 voxel centers, which results in a data set with 2180 grains. For each grain, we took the average over 10 renditions. We varied $R_{n}$ between $4\cdot R_{1}$ and $24\cdot R_{1}$. The results from Figure 3 and 4 refer to the choice $R_{n}=24\cdot R_{1}$.
Figure 3 shows histograms of the estimated volumes and surface areas, respectively. The distributions peak at very small values, but they exhibit comparatively strong tails. Interestingly, the tails are more pronounced for the volumes than the surface areas. We also estimated the tensors $\Phi_{k}^{0,2}$, $k\in\{0,1,2,3\}$. Figure 4 shows a histogram of the ratios between the smallest and biggest absolute eigenvalue for the tensors $\Phi_{1}^{0,2}$ and $\Phi_{2}^{0,2}$. These ratios represent scalar indices of anisotropy, which intuitively quantify the degree of anisotropy with respect to the corresponding Minkowski tensor, and which have already been successfully applied in physics [56, 57, 58, 34, 37]. For the metallic grains, these indices detect a higher degreee of anisotropy with respect to the curvature than the surface area.
6.2 Nanorough surfaces
Figure 5: Examples of nanorough surfaces from Table 10 with three different degrees of roughness, i.e., an RMS value of (a) 7 nm, (b) 24 nm, and (c) 35 nm. The data is taken from [59]; the figures have been created by Jens Uwe Neurohr.
Table 10: Estimated values of the surface and the ratio of the eigenvalues of the tensor $\Phi_{2}^{0,2}$ in comparison with triangulation results based on data from [59]. For the surface estimation with the result of Theorem 4.5 we used Remark 4.6. We used several samples with the same RMS value. The underlying experimental data of the nanorough surfaces is from [59].
Triangulation Least Squares Theorem 4.5
RMS Surface Ratio EV Surface Ratio EV Surface Ratio EV
$[\mathrm{nm}]$ $[\mu\mathrm{m}^{2}]$ $[\mu\mathrm{m}^{2}]$ $[\mu\mathrm{m}^{2}]$
7 11.6 0.285 $13.6$ 0.102 19.1 0.204
11.6 0.287 $13.5$ 0.110 19.1 0.204
11.7 0.295 $13.7$ 0.138 19.1 0.205
24 12.7 0.371 $15.6$ 0.205 19.6 0.231
12.4 0.359 $15.0$ 0.205 19.5 0.227
12.4 0.308 $15.1$ 0.098 19.6 0.229
35 12.2 0.343 $14.7$ 0.160 19.9 0.255
12.5 0.360 $15.2$ 0.184 20.0 0.259
12.3 0.326 $14.9$ 0.155 19.9 0.253
12.4 0.354 $15.0$ 0.178 19.9 0.255
Next, we apply our algorithm to a distinctly different experimental data set, a nanorough surface from [59], measured via atomic force microscopy (AFM). The roughness at the nano scale, i.e., an order of magnitude above atomic resolution, was obtained by a specific etching protocol; see [59] for further details. Both protocol and three-dimensional visualizations of the data sets (see Figure 5) suggest that the nanorough surface can be described (or at least approximated) by a finite union of compact sets with positive reach.
We analyze ten samples in total with three different values of the root mean square (RMS) of height distribution (7 nm, 24 nm, and 35 nm). Each sample is measured within a scan window of size $3\,\mu\text{m}\times 3\,\mu\text{m}$ and a resolution of $512\times 512$ pixels, i.e., each surface is represented by a grid of three-dimensional points, where the $z$-coordinate represents the height of the surface at the corresponding $(x,y)$ position. We varied $R_{n}$ between 12.5 and 250. The results from Table 10 refer to the value $R_{n}=25$, which is about 4 times the average nearest neighbor distance in the data. This rule of thumb balances robustness (due to a large range of radii) and accuracy (resolving fine details of the nanorough structure).
We compare our estimates with those from a triangulated data set as in [59] using methods from [58]. Considering how rough our samples are, and hence challenging for our method of extrapolation, we observe a relatively good agreement of our method with the results from the triangulated data. The Voronoi estimator tends to estimate slightly larger values of the surface area and higher degrees of anisotropy (i.e., smaller values of the anisotropy index). Importantly, the trend between samples agrees for both methods.
Acknowledgments
We thank Karin Jacobs and Christian Spengler for providing their data of nanorough surfaces, and Jens Uwe Neurohr for Figure 5. This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265, under grant numbers HU 1874/5-1, ME 1361/16-1, KL 3391/2-2, WI 5527/1-1, and LO 418/25-1, as well as by the Helmholtz Association and the DLR via the Helmholtz Young Investigator Group “DataMat”.
Appendix A Further simulation results
In this section, we provide additional simulation results for the algorithm described in the paper, based on solving a least-squares problem. The following tables can be found on the subsequent pages:
  • Table 11: Simulation results for the 2-dimensional rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$
  • Table 12: Simulation results for a 2-dimensional rectangle with rounded vertices. Specifically, we mean the parallel set of a rectangle. The parallel set of a compact set $K$ with parameter $r_{0}$ is the set of all points with a distance less than $r_{0}$ to $K$.
  • Table 13: Simulation results for the 3-dimensional rectangle $[-\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$.
Table 11: Results for the 2-dimensional rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=50$ and $R_{n}=2$. We took the average of 50 renditions. All values are rounded to the fourth significant digit.
Tensors Values Algorithm Tensors Values Algorithm
 $\Phi_{0}^{0,0}$ 1 0.9999 $\Phi_{1}^{0,0}$ 8 8.000
[…]
($\Phi_{0}^{2,0}$)2,2 3.125 3.125 ($\Phi_{1}^{2,0}$)1,1 6.75 6.750
[…]
($\Phi_{2}^{2,0}$)2,2 15.63 15.62 ($\Phi_{0}^{1,1}$)1,1 0.4775 0.4774
[…]
Table 12: Results for a 2-dimensional rectangle with rounded vertices intersected with a grid of resolution $a=0.005$. Here $R_{r_{0}}$ denotes the parallel set of the rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ with parameter $r_{0}>0$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 10 renditions.
Tensors Values Algorithm Tensors Values Algorithm
 $\Phi_{0}^{0,0}$ $(R_{\frac{1}{4}})$ 0.999 1 $\Phi_{1}^{0,0}$ $(R_{\frac{1}{4}})$ 8.784 8.785
$\Phi_{2}^{0,0}$ $(R_{\frac{1}{4}})$ 19.194 19.196 $\Phi_{1}^{0,2}$ $(R_{\frac{1}{4}})$ $\begin{pmatrix}0.269&0\\ 0&0.429\end{pmatrix}$ $\begin{pmatrix}0.270&0\\ 0&0.429\end{pmatrix}$
$\Phi_{0}^{0,0}$ $(R_{\frac{1}{2}})$ 0.999 1 $\Phi_{1}^{0,0}$ $(R_{\frac{1}{2}})$ 9.571 9.571
$\Phi_{2}^{0,0}$ $(R_{\frac{1}{2}})$ 23.782 23.785 $\Phi_{1}^{0,2}$ $(R_{\frac{1}{2}})$ $\begin{pmatrix}0.301&0\\ 0&0.460\end{pmatrix}$ $\begin{pmatrix}0.301&0\\ 0&0.460\end{pmatrix}$
$\Phi_{0}^{0,0}$ $(R_{1})$ 0.999 1 $\Phi_{1}^{0,0}$ $(R_{1})$ 11.143 11.142
$\Phi_{2}^{0,0}$ $(R_{1})$ 34.134 34.142 $\Phi_{1}^{0,2}$ $(R_{1})$ $\begin{pmatrix}0.363&0\\ 0&0.522\end{pmatrix}$ $\begin{pmatrix}0.364&0\\ 0&0.523\end{pmatrix}$
Table 13: Results for the 3-dimensional rectangle $[-\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 28 renditions.
Tensors Values Algorithm Tensors Values Algorithm
 $\Phi_{0}^{0,0}$ 1 1.005 $\Phi_{1}^{0,0}$ 6 5.986
$\Phi_{2}^{0,0}$ 11 11.051 ($\Phi_{0}^{0,2}$)1,1 0.080 0.075
($\Phi_{0}^{0,2}$)2,2 0.080 0.079 ($\Phi_{0}^{0,2}$)3,3 0.080 0.078
($\Phi_{1}^{0,2}$)1,1 0.398 0.406 ($\Phi_{1}^{0,2}$)2,2 0.318 0.319
($\Phi_{1}^{0,2}$)3,3 0.239 0.243 ($\Phi_{2}^{0,2}$)1,1 0.478 0.472
($\Phi_{2}^{0,2}$)2,2 0.239 0.238 ($\Phi_{2}^{0,2}$)3,3 0.159 0.156
($\Phi_{0}^{2,0}$)1,1 0.125 0.126 ($\Phi_{0}^{2,0}$)2,2 0.5 0.500
($\Phi_{0}^{2,0}$)3,3 1.125 1.126 ($\Phi_{1}^{2,0}$)1,1 0.667 0.665
($\Phi_{1}^{2,0}$)2,2 2.333 2.333 ($\Phi_{1}^{2,0}$)3,3 4.5 4.497
($\Phi_{2}^{2,0}$)1,1 0.958 0.960 ($\Phi_{2}^{2,0}$)2,2 2.833 2.835
($\Phi_{2}^{2,0}$)3,3 5.625 5.629 ($\Phi_{3}^{2,0}$)1,1 0.25 0.249
($\Phi_{3}^{2,0}$)2,2 1 0.999 ($\Phi_{3}^{2,0}$)3,3 2.25 2.248
($\Phi_{0}^{1,1}$)1,1 0.159 0.159 ($\Phi_{0}^{1,1}$)2,2 0.318 0.319
($\Phi_{0}^{1,1}$)3,3 0.478 0.475 ($\Phi_{1}^{1,1}$)1,1 0.796 0.795
($\Phi_{1}^{1,1}$)2,2 1.273 1.272 ($\Phi_{1}^{1,1}$)3,3 1.432 1.437
($\Phi_{2}^{1,1}$)1,1 0.955 0.958 ($\Phi_{2}^{1,1}$)2,2 0.955 0.956
($\Phi_{2}^{1,1}$)3,3 0.955 0.952